### v1

# IRS on orbits and the case of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$

Abstract.
Let $G$ be a connected Lie group and $\text{Sub}_{G}$ be the space of closed subgroups of $G$ equipped with the Chabauty topology. In this article, we investigate the existence of invariant random subgroups of $G$ supported on various orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. We also find the orbit closure for each of these orbits in the case of the Lie group $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
Key words and phrases:
Connected Lie groups, spaces of closed subgroups, Chabauty topology, invariant random subgroups, conjugacy limits
1991 Mathematics Subject Classification:
Primary: $22$E15151515, $28$D15151515; secondary: $37$A15151515, $37$B02020202
1. Introduction
Let $G$ be a connected (real) Lie group and ${\rm Sub}_{G}$ be the space of (closed) subgroups of $G$ equipped with the Chabauty topology introduced by Claude Chabauty in [4]. The Chabauty topology on SubG is generated by the subsets $\{H\in{\rm Sub}_{G}\mid H\cap K=\emptyset\}$, $\{H\in{\rm Sub}_{G}\mid H\cap U\not=\emptyset\}$, where $K\subset G$ is compact and $U\subset G$ is open. With respect to the Chabauty topology, $\text{Sub}_{G}$ is a compact metrizable space (cf. Lemma E.1.1 of [3]). Moreover, a sequence $H_{n}\in\text{Sub}_{G}$ converges to $H\in\text{Sub}_{G}$ if and only if for any $h\in H$ there is a sequence $h_{n}\in H_{n}$ such that $h_{n}$ converges to $h$ and for any sequence $x_{k_{n}}\in H_{k_{n}}$, with $k_{n+1}>k_{n}$ and $x_{k_{n}}\to x$, we have $x\in H$ (cf. Lemma E.1.2 of [3]).
The group $G$ acts on SubG by conjugation: $(g,H)\mapsto gHg^{-1}$. An invariant random subgroup (hereafter IRS) is a probability measure $\mu$ on SubG that is invariant under the conjugation action of $G$ on SubG. We say that a Borel subset ${\mathcal{O}}$ of SubG supports an IRS $\mu$ if $\mu({{\mathcal{O}}})=1$: a Borel subset ${{\mathcal{O}}}$ supporting an IRS $\mu$ need not be the support of $\mu$ but $\overline{{\mathcal{O}}}$ contains the support of $\mu$. Dirac measure supported on a closed normal subgroup of a connected Lie group is an example of an IRS. Suppose $\Gamma$ is a lattice inside the connected Lie group $G$, i.e., $\Gamma$ is a discrete subgroup of $G$ such that $G/\Gamma$ carries a probability measure $\mu$ that is invariant under the natural action of $G$ on $G/\Gamma$. Then the map from $G$ to $\text{Sub}_{G}$ sending $g$ to $g\Gamma g^{-1}$ factors through $G/\Gamma$, and the push forward of $\mu$ by this map becomes a conjugation-invariant probability measure supported on conjugacy classes of $\Gamma$. So, invariant random subgroups can be thought of as a generalization of both normal subgroups and lattices.
Invariant random subgroups arise naturally as stabilizers of measure preserving actions of $G$ on a probability space $(X,\mu)$. By a result of Varadarajan ([14]), the stabilizer of almost every point of this action is closed in $G$. Moreover, the stabilizer map from $X$ to $\text{Sub}_{G}$ given by $x\to G_{x}$ is measurable. Then, one can push the measure $\mu$ forward to an IRS of $G$. In fact, every IRS can be obtained this way (see [1], [2] for more details).
In recent times, IRS has attracted many studies (see for examples [1] and [2]). In this article, we are interested in the existence of invariant random subgroups supported on certain orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. We start with the orbit of a semisimple Levi subgroup. Let $G$ be a connected Lie group with a closed semisimple Levi subgroup $L$; it may be noted that semisimple Levi subgroups of connected Lie groups need not be closed. But, Levi subgroups of linear groups are closed. Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$. Then $L_{G}=\{xLx^{-1}\mid x\in N\}$ where $N$ is the nilradical of $G$ (cf. Theorem 3.18.3 of [13]). An IRS $\mu$ is called a Levi IRS if $\mu(L_{G})=1$.
Theorem 1.1.
Let $G$ be a connected Lie group with nilradical $N$, and $L$ be a semisimple Levi subgroup of $G$. Then the following are equivalent:
  1. (1)
    there is an IRS supported on the set of all semisimple Levi subgroups;
  2. (2)
    there is a $N$-invariant probability measure supported on semisimple Levi subgroups;
  3. (3)
    the semisimple Levi subgroup $L$ is normal in $G$.
In particular, the set of semisimple Levi subgroups $L_{G}$ is closed in SubG if and only if $L$ is normal.
Corollary 1.2.
Any Levi IRS of a connected Lie group $G$ is a Dirac measure supported on a Levi subgroup of $G$.
So, any connected Lie group $G$ admitting a Levi IRS is of the form $G=L\times S$ where $L$ is semisimple and $S$ is solvable. In particular, $G=\operatorname{SL}(2,\mathbb{R})\times\mathbb{R}^{2}$ admits a Levi IRS, whereas $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ does not admit a Levi IRS.
As in the case of Levi subgroups, the set of all maximal compact subgroups of a connected Lie group is also the orbit of a maximal compact subgroup under the conjugation action. We next provide a necessary and sufficient condition for the set of maximal compact subgroups to support an IRS.
Theorem 1.3.
Let $G$ be a connected Lie group and $K$ be a maximal compact subgroup of $G$. Also, let $K_{G}$ be the set of all maximal compact subgroups of $G$. Then, there is an IRS supported on $K_{G}$ if and only if $K$ is normal in $G$.
We now assume that $G$ is a real algebraic group, that is, $G$ is the group of $\mathbb{R}$-points of an algebraic group defined over $\mathbb{R}$. Let $B$ be a Borel subgroup of $G$, i.e., $B$ is a maximal connected solvable subgroup of $G$. Also, let $B_{G}$ be the set of all Borel subgroups of $G$. Then $B_{G}$ is the orbit of $B$ under the conjugation action. Suppose $H$ is a locally compact group which acts continuously on a locally convex topological vector space $V$. Then, every $H$-invariant, compact, convex subset of $V$ is called a compact, convex $H$-space. We call $H$ to be amenable if $H$ has a fixed point in every compact, convex $H$-space. Abelian, compact and solvable Lie groups are examples of amenable groups. The readers are referred to Chapter $12$ of [12] for more details regarding amenable groups. The next theorem is about the existence of an IRS supported on $B_{G}$.
Theorem 1.4.
Let $G$ be a connected real algebraic group and $B_{G}$ be the set of all Borel subgroups of $G$. Then $B_{G}$ is closed in SubG and there is an IRS supported on $B_{G}$ if and only if $G$ is amenable.
Example 1.5.
Let $G=SO(3,\mathbb{R})\ltimes\mathbb{R}^{3}$. We identify $O(2,\mathbb{R})$ as a subgroup of $SO(3,\mathbb{R})$ by the map $\alpha\mapsto\begin{pmatrix}\alpha&0\cr 0&\det(\alpha)\end{pmatrix}$. Then $B=SO(2,\mathbb{R})\ltimes R^{3}$ is a Borel subgroup of $G$ and the normalizer of $B$ is $O(2,\mathbb{R})\ltimes\mathbb{R}^{3}$. Thus, $SO(3,\mathbb{R})$-invariant measure on the quotient space $SO(3,\mathbb{R})/O(2,\mathbb{R})$ induces an IRS on the set of Borel subgroups $B_{G}$ of $G$.
We next consider the orbit of a maximal diagonalizable subgroup $D$. Here, by diagonalizable subgroup, we mean diagonalizable over $\mathbb{R}$, that is, an $\mathbb{R}$-split torus. The set of all maximal diagonalizable subgroups then consists of conjugates of $D$, we denote this set by $D_{G}$.
Theorem 1.6.
Let $G$ be a connected real algebraic group and $D_{G}$ be the set of all maximal diagnolizable subgroups of $G$. Then there is an IRS supported on $D_{G}$ if and only if $D$ is central in $G$.
We have seen that the group $\operatorname{SL}(2,\mathbb{R})\times\mathbb{R}^{2}$ admits a Levi IRS, whereas, the group $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ does not admit a Levi IRS. In particular, the orbit of a semisimple Levi subgroup is not closed in $\text{Sub}_{G}$ when $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. This leads to a natural question of studying the closure of the orbit under the conjugation action of a semisimple Levi subgroup. The same question can be asked for other orbits discussed above as well. In this article, we address these questions in the case of the Lie group $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Note that, studying these questions is the same as investigating about conjugacy limits of various subgroups of $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. The study of conjugacy limits is an interesting problem in its own right and has been an active area of research aimed at understanding the space $\text{Sub}_{G}$. For example, in [10], [11], the author has studied the conjugacy limits of diagonal subgroups in $\operatorname{SL}(3,\mathbb{R})$ and $\operatorname{SL}(n,\mathbb{R})$ respectively. One may also look at the references cited in these two papers for related works and their relation to geometry.
Given a closed subgroup $H$ of $G$, by a conjugacy limit of $H$ in $\text{Sub}_{G}$, we mean the limit of $(g_{n}Hg_{n}^{-1})$ for a sequence $(g_{n})$ in $G$. Note that $(g_{n}Hg_{n}^{-1})$ always has a convergent subsequence as $\text{Sub}_{G}$ is compact. The following theorem gives a complete description of all the limit points of the set $L_{G}$. We write any element of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ as $(g,v)$ with $g\in\operatorname{SL}(2,\mathbb{R})$ and $v\in\mathbb{R}^{2}$. Also, for any two elements $(g,v)$ and $(g^{\prime},v^{\prime})$, the group operation is defined as follows:
$(g,v)(g^{\prime},v^{\prime})=(gg^{\prime},v+g(v^{\prime})).$
We identify $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ as closed subgroups of $G$ by the maps $g\mapsto(g,(0,0))$ and $v\mapsto(I,v)$ where $I$ denotes the $2\times 2$ identity matrix. For any $c\in\mathbb{R}$, we define a unipotent subgroup $N^{+}_{c}$ by
$N^{+}_{c}=\left\{\big{(}\begin{pmatrix}1&x\\ 0&1\end{pmatrix},(cx,0)\big{)}:x\in\mathbb{R}\right\}.$
For $c=0$, we denote $N^{+}_{c}$ by $N^{+}$, which is identified (and isomorphic) to the group of unipotent upper triangular matrices in $\operatorname{SL}(2,\mathbb{R})$.
Theorem 1.7.
Any subgroup in $\overline{L_{G}}$ is either a Levi subgroup of $G$ or a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
Next, we consider the conjugacy limit of maximal compact subgroups, that is, the closure of $K_{G}=\{gKg^{-1}\mid g\in G\}$ for any maximal compact subgroup $K$ of $G=SL(2,\mathbb{R})\times\mathbb{R}^{2}$. For example, we can take $K=SO(2,\mathbb{R})$. For this purpose, we introduce the following subgroups of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$:
$V_{c}=\{(I,(t,ct))\mid t\in\mathbb{R}\},\ c\in\mathbb{R},~{}~{}V_{\infty}=\{(I,(0,t))\mid t\in\mathbb{R}\}$
and
$\tilde{N}^{+}=\left\{\pm\begin{pmatrix}1&s\\ 0&1\end{pmatrix}:s\in\mathbb{R}\right\}.$
It may be noted that $V_{c},c\in\mathbb{R}\cup\{\infty\}$ is the set of all one-dimensional subspaces of $\mathbb{R}^{2}$, the projective space ${\mathbb{P}}^{1}$.
Theorem 1.8.
The closure of $K_{G}$ consists of conjugates of the subgroups $V_{c},c\in\mathbb{R}\cup\{\infty\}$ and $\tilde{N^{+}}$, in addition to the conjugates of $K$.
The next result identifies limit points of the set of maximal diagonalizable subgroups of $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Let $A=\left\{\begin{pmatrix}a&0\\ 0&\frac{1}{a}\end{pmatrix}\mid a\in\mathbb{R}^{+}\right\}$ and $D_{G}=\{gAg^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}A$. We now introduce the following one-dimensional subgroups
$V_{(a,b,c)}=\left\{\left(\begin{pmatrix}1&at\\ 0&1\end{pmatrix},\left(ct+\frac{abt^{2}}{2},bt\right)\right)\mid t\in\mathbb{R}\right\}$
for any $a,b,c\in\mathbb{R}$. It may be noted that $V_{a,b,c}$’s are the one-parameter subgroups in the $3$-dimensional Heisenberg group and $V_{(1,0,c)}=N^{+}_{c}$, $V_{(0,1,c)}=V_{c}$ and $V_{(0,0,1)}=V_{\infty}$.
Theorem 1.9.
The closure of $D_{G}$ consists of conjugates of $V_{(r_{1},r_{2},r_{3})}$ for any $r_{1},r_{2},r_{3}\in\mathbb{R}$ with $r_{1}=1$ or $r_{3}=1$ or $r_{1}=0$ and $r_{2}=1$ in addition to the conjugates of the diagonal group $A$.
Let $B$ denote the group of all upper triangular matrices in $\operatorname{SL}(2,\mathbb{R})$. Then $B$ is a Borel subgroup of $\operatorname{SL}(2,\mathbb{R})$ and $B\ltimes\mathbb{R}^{2}$ is a Borel subgroup of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Then, using the Iwasawa decomposition in $\operatorname{SL}(2,\mathbb{R})$, it is easy to see that, for conjugacy limits of $B\ltimes\mathbb{R}^{2}$, we only need to consider conjugation by elements of $K=\operatorname{SO}(2,\mathbb{R})$, hence conjugates of $B$ form a closed subset (but by Theorem 1.4 no IRS is supported on the conjugates of $B$ as $G$ is not amenable). We therefore consider conjugacy limits of $B$ inside $G$ and identify all possible conjugacy limits of $B$
Theorem 1.10.
The closure of $\{gBg^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}$ consists of conjugates of $B$, $N^{+}\ltimes\{(t,0)\mid t\in\mathbb{R}\}~{}(\simeq\mathbb{R}^{2})$ and $\mathbb{R}^{2}$.
The next result is concerned about the conjugacy limits of the unipotent subgroup $N^{+}$.
Theorem 1.11.
The closure of $\{gN^{+}g^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}$ consists of conjugates of $N^{+}$ and $V_{0}$.
2. Preliminary
In this section, we prove certain technical results which are needed to discuss the existence or non-existence of IRS on orbits of the conjugation action of $G$ on $\text{Sub}_{G}$. Given a finite-dimensional vector space $V$, let ${\rm Gr}_{k}(V)$ be the be the Grassnannian manifold of $k$-dimensional subspaces of $V$. Equip $V$ with an inner product, and for a subspace $W$ of $V$, let $P_{W}$ denote the orthogonal projection onto the subspace $W$. We define a metric on ${\rm Gr}_{k}(V)$ using the operator norm ||.||||.||| | . | | of $P_{W}$’s as follows: for $W,W^{\prime}\in{\rm Gr}_{k}(V)$,
$d(W,W^{\prime})=||P_{W}-P_{W^{\prime}}||.$
The following Lemma is a comparison between the Grassmannian topology and the Chabauty topology of ${\rm Gr}_{k}(V)$ seen as a subset of $\text{Sub}_{V}$.
Lemma 2.1.
Let $V$ be a finite-dimensional vector space and ${\rm Gr}_{k}(V)$ be the Grassmannian manifold of $k$-dimensional subspaces of $V$. Let $\mathfrak{i}\colon{\rm Gr}_{k}(V)\to{\rm Sub}_{V}$ be the canonical inclusion. Then ${\rm Gr}_{k}(V)$ is homeomorphic to its image. In particular, the set of all subspaces of dimension $k$ is a closed set in Chaubauty topology.
Proof.
Suppose $W_{n}\to W$ in ${\rm Gr}_{k}(V)$. This implies that $||P_{W_{n}}-P_{W}||\rightarrow 0$ as $n\rightarrow\infty$. Now let $w\in W$, we want to show that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$. Let $w_{n}=P_{W_{n}}(w)$. Then clearly $w_{n}\in W_{n}$ and
$||w_{n}-w||=||P_{W_{n}}(w)-P_{W}(w)||\leq||P_{W_{n}}-P_{W}||||w||\rightarrow 0$
as $n\rightarrow\infty$. On the other hand assume that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$, we want to show that $w\in W$. It is enough to show that $P_{W}(w)=w$. Since $w_{n}\rightarrow w$ and $P_{W_{n}}\rightarrow P_{W}$, it follows that $P_{W_{n}}(w_{n})\rightarrow P_{W}(w)$. But as $P_{W_{n}}(w_{n})=w_{n}$, and $w_{n}\rightarrow w$, it follows that $P_{W}(w)=w$. Thus, $W_{n}\to W$ in SubV. This proves that the map $\pi$ is continuous. Since ${\rm Gr}_{k}(V)$ is compact, ${\rm Gr}_{k}(V)$ is homeomorphic to $\mathfrak{i}({\rm Gr}_{k}(V))$. ∎
It is known that if a sequence of closed subgroups $H_{n}$ of a Lie group $G$ converges to a closed subgroup $H$, then dim$(H)\geq{\rm limsup}~{}\text{dim}(H_{n})$ (cf. Theorem 1 of [15]). However, in the case of closed subgroups of thev same dimension and in the case of exponential Lie groups, we have the following.
Proposition 2.2.
Let $G$ be a connected Lie group and $H_{n}\to H$ in SubG. Let $\mathcal{G}$, $\mathcal{H}_{n}$ and $\mathcal{H}$ be Lie algebras of $G$, $H_{n}$ and $H$ respectively.
  1. (1)
    If $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$, then $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
  2. (2)
    If $G$ is an exponential Lie group and $H_{n}$’s are connected, then $H$ is also connected, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG and ${\rm limsup}~{}\text{dim}(H_{n})=\text{dim}(H)$.
Remark 2.3.
  1. (1)
    The map sending a closed subgroup to its Lie algebra is known to be a measurable map (cf. proof of Lemma 6 in [16]) but here we prove continuity under some restrictions though measurability is sufficient for our purpose.
  2. (2)
    It may be noted that $H$ need not always be connected even when $H_{n}$ are all connected and have the same dimension: e.g., cases $1$ and $2$ of Section $5.3$ provide a sequence of conjugate subgroups whose limit point has two connected components.
Proof.
Let $V$ be a limit of $\mathcal{H}_{n}$ in SubG. By passing to a subsequence, we may assume that $\mathcal{H}_{n}\to V$ in SubG. Then $V$ is a vector subspace. In is easy to see that it is a Lie subalgebra of $\mathcal{G}$ as well. For $v\in V$, there are $v_{n}\in\mathcal{H}_{n}$ such that $v_{n}\to v$, hence $tv_{n}\to tv$ for all $t\in\mathbb{R}$. Let $\exp\colon\mathcal{G}\to G$ be the exponential map. Then $\exp(tv_{n})\to\exp(tv)$ for all $t\in\mathbb{R}$. Since $H_{n}$ converges to $H$, $\exp(tv)\in H$ for all $t\in\mathbb{R}$. This implies that $v\in\mathcal{H}$. Thus, $V\subset\mathcal{H}$.
Assume that $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$. Then $\text{dim}(\mathcal{H}_{n})=\text{dim}(\mathcal{H})$ for all $n$. By Lemma 2.1, $\text{dim}(V)=\text{dim}(\mathcal{H})$. Since $V\subset\mathcal{H}$, we get that $V=\mathcal{H}$. Thus, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
Assume that $\exp\colon\mathcal{G}\to G$ is a diffeomorphism and $H_{n}$ are connected. Then $\exp\colon\mathcal{H}_{n}\to H$ is also a diffeomorphism for all $n$. Let $h\in H$. Since $H_{n}\to H$, there are $h_{n}\in H_{n}$ such that $h_{n}\to h$. Let $X_{n}\in\mathcal{H}_{n}$ and $X\in\mathcal{H}$ be such that $\exp(X_{n})=h_{n}$ and $\exp(X)=h$. Since $\exp$ is a diffeomorphism, we get that $X_{n}\to X$. Since $\mathcal{H}_{n}\to V\subset\mathcal{H}$, we get that $X\in\mathcal{H}$. Thus, $\exp(\mathcal{H})=H$. Therefore, $H$ is connected. Since $\exp$ is a diffeomorphism, we also get that $\mathcal{H}_{n}\to\mathcal{H}$ and hence $V=\mathcal{H}$. Thus, ${\rm limsup}~{}\text{dim}(H_{n})=\text{dim}(H)$.
The result of the following kind is known as Furstenberg’s Lemma (cf. Section 5, [7]) and it has been proved for minimally almost periodic groups in Lemma 3 of [6] and when $\Gamma$ (as in Lemma 2.4) is certain semisimple analytic group in Proposition 5.1 of [7]. Here we provide the following using a result of [5].
Lemma 2.4.
Let $V$ be a finitedimensional vector space and $\Gamma$ be a subgroup of $GL(V)$. Also, let $U$ be a subgroup of $\Gamma$ consisting of unipotent transformations. Let ${\rm Gr}_{k}(V)$ be the Grassmannian manifold of $k$-dimensional subspaces of $V$ and $\lambda$ be a probability measure on ${\rm Gr}_{k}(V)$ which is invariant under the induced action of $\Gamma$ on ${\rm Gr}_{k}(V)$. Also, let $I_{\lambda}=\{\tau\in GL(V)\mid\tau~{}~{}{\rm is~{}~{}trivial~{}~{}on~{}}S\}$ where $S$ is the support of $\lambda$. Then $\Gamma$ is contained in an algebraic group $\tilde{\Gamma}$ such that $\tilde{\Gamma}/I_{\lambda}$ is a compact group, and $S$ consists of $U$-invariant subspaces.
Proof.
Let $\lambda$ be the $\Gamma$-invariant probability measure on ${\rm Gr}_{k}(V)$. Let $G_{\lambda}=\{\tau\in GL(V)\mid\tau(\lambda)=\lambda\}$ and $I_{\lambda}=\{\tau\in GL(V)\mid\tau~{}~{}{\rm is~{}~{}trivial~{}~{}on~{}}S\}$ where $S$ is the support of $\lambda$. Then $I_{\lambda}$ is a normal algebraic subgroup of $G_{\lambda}$. Since ${\rm Gr}_{k}(V)\simeq GL(V)/H$ where $H$ is the stablizer of a $k$-dimensional subspace, by Corollary 2.6 of [5] we get that $G_{\lambda}$ is a real algebraic group and $G_{\lambda}/I_{\lambda}$ is compact. Thus, any unipotent transformation in $G_{\lambda}$ is in $I_{\lambda}$. This proves the result. ∎
Proposition 2.5.
Let $G$ be a connected Lie group with a closed connected Lie subgroup $S$. Let $\Gamma$ be a $\sigma$-compact subgroup of Aut$(G)$ and $U$ be a subgroup of $\Gamma$ consisting of unipotent automorphisms. Suppose there is a ${\Gamma}$-invariant probability measure $\mu$ on ${\rm Sub}_{G}$ such that $\mu(\{\tau(S)\mid\tau\in\Gamma\})=1$. Then $S$ is $N$-invariant where $N$ is the smallest normal subgroup of $\Gamma$ containing $U$.
Proof.
Let $\mathcal{G}$ and $\mathcal{S}$ be the Lie algebras of $G$ and $S$ respectively. Let $\mathcal{O}=\{\tau(S)\mid\tau\in{\rm Aut}~{}(G)\}$. Let $X={\rm Gr}_{k}(\mathcal{G})$ be the Grassmannian manifold of $k$-dimensional subspaces of $\mathcal{G}$ where $k$ is the dimension of $\mathcal{S}$. Let $\phi\colon\mathcal{O}\to X$ be defined by $\phi(\tau(S))={\rm d}\tau(\mathcal{S})$. Then it follows from Proposition 2.2 and Lemma 2.1 that $\phi$ is continuous.
Let $\lambda=\phi(\mu)$. Consider the canonical action of $\Gamma$ on $X$ given by: $(\tau,V)\mapsto{\rm d}\tau(V)$ for $\tau\in\Gamma$ and $V\in X$. Then $\lambda$ is a $\Gamma$-invariant probability measure on $X$. By Lemma 2.4, $\lambda$ is supported on $U$-invariant subspaces. This proves that $\mathcal{S}$ is $N$-invariant. Since $S$ is a connected Lie subgroup, $S$ is $N$-nvariant. ∎
3. IRS on orbits
In this section, we show the existence and non-existence of invariant random subgroups supported on various orbits of the conjugation action of a connected Lie group $G$ on SubG. In other words, we discuss the proofs of Theorems 1.1, 1.3, 1.4 and 1.6. We also discuss certain structural results related to these theorems.
Proof of Theorem 1.1.
It is sufficient to prove that (2) implies (3). Let $L$ be a semisimple Levi subgroup of $G$ and $N$ be the nilpotent radical of $G$ as in the statement of the theorem. Then $L$ is a connected Lie subgroup of $G$. Suppose that there is an IRS supported on the set $L_{G}$ of all semisimple Levi subgroups. Since any two semisimple Levi subgroups conjugate in $G$ by an element of $N$, there is a $N$-invariant probability measure $\mu$ such that $\mu(\{xLx^{-1}\mid x\in N\})=1$. Let $\mathcal{G}$ be the Lie algebra of $G$. Since $N$ is a nilpotent normal subgroup of $G$, Ad$(x)$ is a unipotent transformation on $\mathcal{G}$ for all $x\in N$. Therefore, by Proposition 2.5, $\mu$ is supported on $N$-invariant subspaces of $\mathcal{G}$. Thus, $L$ is normalized by $N$. Since any two semisimple Levi subgroups are conjugate by an element of $N$, we get that $L$ is normal in $G$.
Suppose the set of all semisimple Levi subgroups is closed in SubG. Then $L_{G}$ is compact; and, since $N$ is amenable, $L_{G}$ admits an $N$-invariant probability measure. Therefore, $(2)$ implies $(3)$ proves that $L$ is normal in $G$.
Proof of Theorem 1.3.
Suppose that there is an IRS $\mu$ such that $\mu(K_{G})=1$, where $K_{G}$ is the set of all maximal compact subgroups of $G$ with $K$ being a maximal compact subgroup of $G$. Then $K$ is connected and $K_{G}=\{xKx^{-1}\mid x\in G\}$ (cf. Theorem 3.1, Chapter XV of [9]), hence by Proposition 2.5, $K$ is normalized by any Ad-unipotent subgroup of $G$.
Let $L$ be a semisimple Levi subgroup of $G$ and $K_{l}$ be a maximal compact subgroup of $L$. Then $K_{l}\subset gKg^{-1}$ for some $g\in G$. Replacing $K$ by a conjugate, we may assume that $K$ contains $K_{l}$. This, in particular, implies that $K$ is normalized by compact simple factors of $L$. Let $L_{1}$ be a noncompact simple factor of $L$. Then $L_{1}$ is generated by the Ad-unipotent elements of $L_{1}$. Therefore, $L_{1}$ normalizes $K$. Thus, the seimsimple Levi subgroup $L$ normalizes $K$.
Now, let $N$ be the nilpotent radical of $G$. Then Ad$(x)$ is unipotent for any $x\in N$, hence $N$ normalizes $K$. Thus, $NK$ is a closed connected subgroup of $G$. Since $K$ is a maximal compact subgroup of $G$, $K$ is a maximal compact subgroup of $NK$. Since $N$ normalizes $K$, $K$ is normal in $NK$ and hence contains any compact subgroup of $NK$. Recall that any two maximal compact subgroups of a connected Lie groups are conjugate. Let $T=N\cap K$. Then $T$ is the largest compact normal subgroup of $N$ and hence $N/T$ is a simply connected nilpotent Lie group. Since $N$ is normal in $G$ and $T$ is the largest compact normal subgroup of $N$, $T$ is normal in $G$. Since $T\subset K$, replacing $G$ by $G/T$, we may assume that $N$ is simply connected.
Let $\tilde{G}=G/N$ and $\tilde{K}$ be a maximal compact subgroup of $\tilde{G}$ containing $KN/N$. Then $\tilde{K}$ is connected and there is a closed subgroup $H$ of $G$ containing $N$ such that $H/N=\tilde{K}$. Since $N$ and $\tilde{K}$ are connected, $H$ is also connected (cf. Theorem 7.4 of [8]). As $N$ is simply connected, $H$ contains a compact subgroup $M$ such that $H=MN$ and $\tilde{K}=H/N\simeq M$ (cf. Theorem 3.2, Chapter XII of [9]). Also, $K\subset H$ as $H/N=\tilde{K}$ contains $KN/N$. Since $M$ is a compact subgroup of $H$ and $K$ is a maximal compact subgroup, $M\subset xKx^{-1}$ for some $x\in H=MN$. But $N$ normalizes $K$, hence $M\subset K$. This shows that $\tilde{K}=KN/N$.
Now, let $R$ be the solvable radical of $G$. Then $N\subset R$ and by Corollary 3.16.4 of [13], $R/N$ is the center of $\tilde{G}$, hence $\tilde{K}$ is normalized by $R/N$. This shows that $KN$ is normalized by $R$. Since the semisimple Levi subgroup $L$ normalizes $K$, $L$ normalizes $KN$ as well. It follows that $KN$ is a normal subgroup of $G$.
As $N$ normalizes $K$, $K$ is a normal subgroup of $KN$. Since $N$ is a simply connected nilpotent Lie group, $N\cap K=\{e\}$ and hence $KN/K~{}(\simeq N)$ has no compact subgroup. This proves that $K$ is a normal subgroup of $G$. ∎
The existence of an IRS supported on the set of maximal compact subgroups has certain implications on the structure of the connected Lie group $G$ under consideration as described in the following proposition.
Proposition 3.1.
Let $G$ be a connected Lie group with solvable radical $R$ and $M$ be a maximal compact subgroup of $G$. Suppose $M$ is normal in $G$. Then we have the following:
  1. (1)
    solvable radical $R$ admits a compact central subgroup $T$ so that $R/T$ is simply connected;
  2. (2)
    any closed simple Lie subgroup of $G$ is either compact or has no nontrivial compact connected subgroups.
Remark 3.2.
The universal cover of $SL(2,\mathbb{R})$ is a simple Lie group without any compact subgroup of positive dimension. Thus the second possibility in 2 of Proposition 3.1 may occur.
Proof.
Suppose the maximal compact subgroup $M$ of $G$ is normal in $G$. If $R$ is the solvable radical of $G$, then $R\cap M$ is a maximal compact subgroup of $R$. Since $M$ is normal in $G$, $R\cap M$ is normal in $R$. If $T=R\cap M$, then $T$ is abelian and $R/T$ has no compact subgroup. Thus, $R/T$ is simply connected (cf. exercise 3, Chapter XII of [9]).
Let $S$ be any simple Lie subgroup of $G$. Suppose that there exists a non-trivial compact connected subgroup of $S$. Then $M\cap S$ is a compact normal subgroup of positive dimension in $S$. Since $S$ is simple, $S=S\cap M$. ∎
Now, we assume that $G$ is a real algebraic group. Also, let $B$ be a Borel subgroup of $G$ and $B_{G}$ be the set of all Borel subgroups of $G$, i.e., $B_{G}=\{xBx^{-1}|x\in G\}$.
Proof of Theorem 1.4.
Since $B$ is a Borel subgroup of $G$, $G/B$ is compact, and hence there is a compact set $C$ in $G$ such that $G=CB$. Since any two Borel subgroups are conjugate in $G$, $B_{G}=\{xBx^{-1}\mid x\in C\}$. Thus, $B_{G}$ is a compact subset of SubG.
Suppose that there is an IRS $\mu$ such that $\mu(B_{G})=1$. To prove that $G$ is amenable, it is sufficient to prove that any semisimple Levi subgroup of $G$ is compact. Let $R$ be the solvable radical of $G$. Then $R\subset B$, hence any Borel subgroup of $G$ is a lift of a Borel subgroup of $G/R$. Thus, replacing $G$ by $G/R$, we may assume that $G$ is semisimple. Now, let $S$ be a simple factor of $G$. If $S$ is non-compact, then $S$ is generated by unipotent elements. This implies by Proposition 2.5 that $B$ is normalized by $S$, hence $B\cap S$ is a solvable normal subgroup of $S$ containing a Borel subgroup of $S$. This is a contradiction to the simplicity of $S$. Hence, any simple factor of $G$ is compact. This completes the proof. ∎
Proof of Theorem 1.6.
Let $D$ be a maximal diagnolizable subgroup of $G$ and $U$ be the unipotent radical of $G$. Then there is a reductive Levi subgroup $L$ of $G$ containing $D$ such that $G=LU$. Let $L^{\prime}=[L,L]$. Then $L^{\prime}$ is a semisimple Levi subgroup of $G$.
Suppose that there is an IRS $\mu$ such that $\mu(D_{G})=1$. Then, by Proposition 2.5, $D$ is normalized by any unipotent subgroup of $G$. If $S$ is a noncompact simple algebraic subgroup of $G$, then maximal split diagonalizable subgroup $D_{S}$ of $S$ is contained in a conjugate of $D$. Replacing $S$ by a conjugate of $S$, we may assume that $D_{S}\subset D$. Since $S$ is noncompact, $S$ is generated by unipotent elements, hence $S$ normalizes $D$. Therefore, $D\cap S$ is a normal subgroup of positive dimension which is a contradiction to the simplicity of $S$. Thus, the semisimple Levi subgroup $L^{\prime}=[L,L]$ of $G$ is compact. This implies that $D\subset Z(L)$ because $L=[L,L]Z(L)$.
Since $D$ is normalized by the unipotent subgroups, $D$ is normalized by $U$. Since $D\cap U$ is trivial, $D$ is centralized by $U$ also. Since $G=LU$, $D$ is central in $G$.
The maximal diagonalizable subgroup being normal has certain implications on the structure of $G$ which is noted down below.
Proposition 3.3.
Let $G$ be a connected real algebraic group and $D$ be a maximal diagonalizable subgroup of $G$. In addition, suppose $D$ is normal in $G$. Then any Levi subgroup of $G$ is a compact subgroup, and the solvable radical is a compact extension of a nilpotent group.
Proof.
Let $H$ be a noncompact simple factor of a semisimple Levi subgroup of $G$ and $D_{H}$ be the maximal diagonalizable subgroup of $H$. Then $D_{H}$ is contained in a conjugate of $D$. Since $D$ is normal, $D_{H}\subset D$. This implies that $D\cap H$ is an abelian normal subgroup of positive dimension in $H$, which is a contradiction to the simplicity of $H$. Therefore, any semisimple Levi subgroup of $G$ is compact.
Now, let $R$ be the solvable radical of $G$. Then $R\cap D$ is a normal subgroup $R$, hence $D_{R}U$ is a nilpotent subgroup of $R$ where $D_{R}=D\cap R$ and $U$ is the unipotent radical of $G$. Since $R/D_{R}U$ is a anisotropic torus (that is, without any split torus subgroup) we get that $R/D_{R}U$ is compact. ∎
4. Application to measure preserving actions
Let $G$ be a locally compact group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations. It follows from [14] that the stabilizer of almost every point in X is closed in G and the stabilizer map $x\mapsto G_{x}$ is measurable where $G_{x}=\{g\in G\mid gx=x\}$ for $x\in X$. Thus, the $G$-invariant probability measure $m$ on $X$ may be realized as an IRS on G which leads to the following:
Corollary 4.1.
Let $G$ be a connected Lie group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations. Then almost none of the stabilizer subgroups is a semisimple Levi subgroup (resp. maximal compact subgroup) of $G$ or $m$ has an atom $x$ such that $G_{x}$ is a semisimple Levi subgroup (resp. maximal compact subgroup) of $G$ and $G_{x}$ is normal in $G$. In addition, if the action of $G$ is ergodic, then almost none of the stabilizer is a semisimple Levi subgroup of $G$ or the action of $G$ factors through the solvable radical of $G$.
Proof.
Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$ and $X_{L}=\{x\in X\mid G_{x}\in L_{G}\}$. Since $L_{G}$ is invariant under the conjugation action and $G_{gx}=gG_{x}g^{-1}$, it follows that $X_{L}$ is invariant under the action of $G$. Suppose $m(X_{L})>0$. Then, replacing $X$ by $X_{L}$ and $m$ by the restriction of $m$ to $X_{L}$, we obtain a measure preserving action of $G$ such that almost all stabilizers are Levi subgroups of $G$. Using the identification of $x\mapsto G_{x}$, we obtain a IRS supported on the set of semisimple Levi subgroups. By Theorem 1.1, we get that $G$ has a unique semisimple Levi subgroup, say L, and $L$ is normal in $G$. The case of maximal compact subgroups may be proved in a similar way.
Suppose the action of $G$ is ergodic. Then $X_{L}$ has full measure or null. In case $X_{L}$ has full measure, any $G_{x}=G_{y}$ for all $x,y\in X_{L}$ and hence the semisimple Levi subgroup acts trivially on $X$, hence the action factors through the solvable radical of $G$.
A similar application can also be done for maximal diagonalizable groups and Borel subgroups as in the following: we omit the details as it is same as in Corollary 4.1
Corollary 4.2.
Let $G$ be a connected real algebraic group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations. Then almost none of the stabilizer subgroup is a Borel (resp. maximal diagonalizable) subgroup of $G$ or $m$ has an atom $x$ such that $G_{x}$ is a Borel (resp. maximal diagonalizable) subgroup of $G$ and $G_{x}$ is normal in $G$. In addition if the action of $G$ is ergodic, then almost none of the stabilizer is a Borel (resp. maximal diagonalizable) subgroup of $G$ or the action of $G$ factors through a compact subgroup (resp. compact extension of the unipotent radical) of $G$.
Corollary 4.3.
Let $G$ be a connected Lie group and $K$ be a maximal compact subgroup of $G$. Suppose $G/K$ carries a $G$-invariant probability measure. Then $G$ is a compact group.
Proof.
Let $m$ be the $G$-invariant probability measure on $G/K$. Then $(G/K,m)$ is a probability space on which $G$ acts by measure preserving transformations and stabilizer at any $xK\in G/K$ is $xKx^{-1}$. Therefore by Corollary 4.1, $K$ is normal in $G$, hence $G/K$ is a locally compact group. Since $G/K$ carries a $G$-invariant probability measure, $G/K$ is compact. Thus, $G$ is compact. ∎
Corollary 4.4.
Let $G$ be a connected real algebraic group and $D$ be a maximal diagonalizable subgroup of $G$. Suppose $G/D$ carries a $G$-invariant probability measure. Then $G$ is a direct product of $D$ and a compact group.
Proof.
Let $m$ be the $G$-invariant probability measure on $G/D$. Then $(G/D,m)$ is a probability space on which $G$ acts by measure preserving transformations and stabilizer at any $xD\in G/D$ is $xDx^{-1}$. Therefore by Corollary 4.2, $D$ is normal in $G$. By Proposition 3.3, the Levi subgroup of $G$ is a compact normal subgroup and the solvable radical is nilpotent. Since $G/D$ has a $G$-invariant probability measure, unipotent radical of $G$ is trivial. Thus, $G$ is a direct product of $D$ and a compact group. ∎
5. Conjugacy limits in $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$
5.1. Limits of levi subgroups
From now onwards, we use the notation $G$ for $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$, and the notation $L$ for the group $\operatorname{SL}(2,\mathbb{R})$ inside $G$. Also, let $L_{G}$ be the set of all semsimple Levi subgroups inside $G$. Then $L_{G}$ consists of conjugates of $L$ by elements of $G$. We denote elements in $G$ by $(g,v)$ with $g\in\operatorname{SL}(2,\mathbb{R})$ and $v\in\mathbb{R}^{2}$. Now, if $(G_{n})$ is a sequence of Levi subgroups in $L_{G}$, then each $G_{n}$ is given by a conjugate of $L$ by some elements of $G$, i.e.,
$G_{n}=\ (g_{n},v_{n})\ L\ (g_{n},v_{n})^{-1},$
where $(g_{n},v_{n})^{-1}$ denotes the inverse of $(g_{n},v_{n})$ in $G$.
If we conjugate $L$ by elements of the form $(g,(0,0))$, the resultant is $L$ itself. Also, note that $(g,v)=(I,v)\ (g,(0,0))$, where $I$ is the identity element in $\operatorname{SL}(2,\mathbb{R})$. Therefore, to determine all the limit points of $L_{G}$, it is enough to consider the conjugacy limits of the form
$(I,v_{n})\ L\ (I,v_{n})^{-1}.$
The following lemma is a key ingredient for the results in this section and afterwards.
Lemma 5.1.
Let $(g_{n})$ and $(v_{n})$ be sequences in $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ respectively such that $v_{n}\not\to 0$ and
$(I,v_{n})\ (g_{n},(0,0))\ (I,v_{n})^{-1}=(g_{n},v_{n}-g_{n}(v_{n}))\rightarrow(g,(0,0))$
for some $g\in\operatorname{SL}(2,\mathbb{R})$. Then $g$ is unipotent.
Proof.
Suppose $(g_{n})$ and $(v_{n})$ are sequences in $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ respectively such that $v_{n}\not\to 0$ and
$(I,v_{n})\ (g_{n},(0,0))\ (I,v_{n})^{-1}=(g_{n},v_{n}-g_{n}(v_{n}))\rightarrow(g,(0,0))$
for some $g\in\operatorname{SL}(2,\mathbb{R})$. Then $g_{n}\to g$ in $SL_{2}(\mathbb{R})$ and $(I-g_{n})(v_{n})\rightarrow(0,0)$ as $n\rightarrow\infty$. Note that if $I-g$ is invertible, then since $g_{n}\to g$, $I-g_{n}\to I-g$. Since invertible matrices form an open set, $I-g_{n}$ is invertible for all large $n$ and hence $(I-g_{n})^{-1}\to(I-g)^{-1}$. As $(I-g_{n})(v_{n})\rightarrow(0,0)$, we arrive at a contradiction to the assumption that $v_{n}\not\to 0$. Therefore, $I-g$ is not invertible. Since $g\in SL(2,\mathbb{R})$, $g$ is an unipotent transformation. ∎
Proof of Theorem 1.7.
Assume that $(I,v_{n})\ L\ (I,v_{n})^{-1}$ converges to $H$ in $\text{Sub}_{G}$; we want to determine $H$. The proof is divided into following caes.
Case $1$. Suppose $v_{n}\rightarrow v$ as $n\rightarrow\infty$. In this case, it is easy to see that $(I,v_{n})\ L\ (I,v_{n})^{-1}$ converges to the closed subgroup $(I,v)\ L\ (I,v)^{-1}$, and hence the limit $H$ is a Levi subgroup.
We now assume that $(v_{n})$ is unbounded. It follows from Lemma 5.1 that if $g\in SL(2,\mathbb{R})\cap H$, then $g$ is unipotent.
Case $2$. Suppose $v_{n}=(\alpha_{n},\beta_{n})$ and $\alpha_{n}$ and $\beta_{n}$ both diverge. We first show that the subgroup $\mathbb{R}^{2}$ is contained in $H$. Let $\mathfrak{g}_{n}=\begin{pmatrix}1&\frac{-s}{\beta_{n}}\\ 0&1\end{pmatrix}$. Then
$(I,v_{n})\ (\mathfrak{g}_{n},(0,0))\ (I,v_{n})^{-1}=(\mathfrak{g}_{n},v_{n}-\mathfrak{g}_{n}(v_{n}))=(\mathfrak{g}_{n},(s,0))\rightarrow(I,(s,0))$
as $n\rightarrow\infty$. Similarly, considering
$\mathfrak{h}_{n}=\begin{pmatrix}1&0\\ \frac{-t}{\alpha_{n}}&1\end{pmatrix}$
it is easy to see that
$(I,v_{n})\ (\mathfrak{h}_{n},(0,0))\ (I,v_{n})^{-1}\rightarrow(0,t)$
as $n\rightarrow\infty$. Then composing $(I,(s,0))$ and $(I,(0,t))$, we see that $(I,(s,t))\in H$ for any $s,t\in\mathbb{R}$.
Since $G=SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ and $\mathbb{R}^{2}\subset H$, we get that $H=(H\cap L)\ltimes\mathbb{R}^{2}$. But $L\cap H$ consists of unipotent transformations. Since $N^{+}$ and its conjugates in $SL(2,\mathbb{R})$ are the maximal subgroups with unipotent elements, $H$ is given by
$H=hN^{+}h^{-1}\ltimes\mathbb{R}^{2}$
for some $h\in\operatorname{SL}(2,\mathbb{R})$, where
$N^{+}=\left\{\begin{pmatrix}1&s\\ 0&1\end{pmatrix}:s\in\mathbb{R}\right\},$
Case $3$. The sequence $(v_{n})=(\alpha_{n},\beta_{n})$ diverges but one of $\alpha_{n}$ or $\beta_{n}$ is bounded. First suppose, $v_{n}=(\alpha_{n},0)$ with $\alpha_{n}$ diverging. Then, for any real number $s$, if we take $\mathfrak{g}_{n}:=\begin{pmatrix}\frac{\displaystyle{\alpha_{n}-s}}{\displaystyle{\alpha_{n}}}&0\\ 0&\frac{\displaystyle{\alpha_{n}}}{\displaystyle{\alpha_{n}-s}}\end{pmatrix}$, we see that
$(I,v_{n})(\mathfrak{g}_{n},(0,0))(I,v_{n})^{-1}\to(I,(s,0)).$
Similarly, considering $\mathfrak{h}_{n}:=\begin{pmatrix}1&0\\ \frac{\displaystyle{-s}}{\displaystyle{\alpha_{n}}}&1\end{pmatrix}$, we see that
$(I,v_{n})(\mathfrak{h}_{n},(0,0))(I,v_{n})^{-1}\to(0,s).$
This shows that $\mathbb{R}^{2}$ is contained in the limit subgroup $H$. Also, since any element of $N^{+}$ fixes $v_{n}$, it follows that $N^{+}$ is contained in $H$ as well. Now it follows, as in Case $2$, that
$H=N^{+}\ltimes\mathbb{R}^{2}.$
Similarly, if $v_{n}=(0,\alpha_{n})$ with $\alpha_{n}$ diverging, then
$H=N^{-}\ltimes\mathbb{R}^{2}.$
where
$N^{-}=\left\{\begin{pmatrix}1&0\\ s&1\end{pmatrix}:s\in\mathbb{R}\right\}.$
Now suppose $v_{n}=(\alpha_{n},\beta_{n})$ with $\alpha_{n}$ diverging and $\beta_{n}$ bounded. As we are looking for limit points of the set of all Levi subgroups, we may assume in this case that $\alpha_{n}$ diverges and $\beta_{n}$ converges to some $\beta$. We can write $(I,(\alpha_{n},\beta_{n}))=(I,(\alpha_{n},0))(I,(0,\beta_{n}))$. So, if $(I,(\alpha_{n},0))L(I,(\alpha_{n},0))^{-1}$ converges to the subgroup $H$, then
$(I,v_{n})L(I,v_{n})^{-1}$ converges to the subgroup $(I,(0,\beta))H(I,(0,\beta))^{-1}$.
Therefore, $H$ is a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
Thus, any subgroup in $\overline{L_{G}}$ is either a Levi subgroup or a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
5.2. Limits of maximal compact subgroups
We now identify the limit points of the set of maximal compact subgroups. In other words, we find the closure of the orbit of $K$ under the conjugacy action of $G$ inside $\text{Sub}_{G}$.
Proof of Theorem 1.8.
Let us recall that $K$ is given by
$K:=\ \left\{\begin{pmatrix}\text{cos}\ \theta&\text{sin}\ \theta\\ -\text{sin}\ \theta&\text{cos}\ \theta\end{pmatrix}\ :\ \theta\in\mathbb{R}\right\}.$
In order to find the limit points of
$K_{G}=\{(g,v)K(g,v)^{-1}\mid g\in SL(2,\mathbb{R}),v\in\mathbb{R}^{2}\},$
it is sufficient to identify the limit points up to conjugation. Using Iwasawa decomposition, i.e., $SL(2,\mathbb{R})=N^{+}AK$, we may assume that $g\in N^{+}A$. Let $H$ be a limit point of $K_{G}$. Then there exist sequences
$\mathfrak{a}_{n}=\begin{pmatrix}a_{n}&0\\ 0&a_{n}^{-1}\end{pmatrix}\in A,~{}\mathfrak{s}_{n}=\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix}\in N^{+}$
and $v_{n}=(\alpha_{n},\beta_{n})\in\mathbb{R}^{2}$ such that $(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})K(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}\to H$. By passing to a subsequence and replacing $\mathfrak{a}_{n}$ or $\mathfrak{s}_{n}$ by their conjugates, we may assume that $a_{n}\to\infty$ or $a_{n}=1$ and $s_{n}\to\infty$ or $s_{n}=0$.
Let $(g,v)\in H$. Then there exists
$\Theta_{n}=\begin{pmatrix}\cos\theta_{n}&\sin\theta_{n}\\ -\sin\theta_{n}&\cos\theta_{n}\end{pmatrix}\in K$
such that $(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})\Theta_{n}(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}\to(g,v)$
Now,
$(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})\Theta_{n}(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}=(T_{n},(x_{n},y_{n})),$
where
$T_{n}=\begin{pmatrix}\cos\theta_{n}-\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}&(\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\\ -\sin\theta_{n}/a_{n}^{2}&\cos\theta_{n}+\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\end{pmatrix}$
and
$x_{n}=\alpha_{n}\big{(}1-\cos\theta_{n}+{\frac{s_{n}}{a_{n}^{2}}}\sin\theta_{n}\big{)}+\beta_{n}\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n},$
$y_{n}=-\alpha_{n}\frac{\sin\theta_{n}}{a_{n}^{2}}+\beta_{n}\big{(}1-\cos\theta_{n}-\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\big{)}.$
We now assume that $v_{n}\to\infty$. As $\big{(}T_{n},(x_{n},y_{n})\big{)}\to(g,v)$, it follows from Lemma 5.1 that $g$ is unipotent.
Case $1$: We now consider the case when ${\frac{\alpha_{n}}{\beta_{n}}}\to 0$. Since $v_{n}$ diverges, $\beta_{n}\to\infty$. Let $x,y\in\mathbb{R}$ be such that $v=(x,y)$. Then
$x_{n}=\alpha_{n}\big{(}1-\cos\theta_{n}+\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\big{)}+\beta_{n}\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n}\to x.$
Since $T_{n}\to g$, $\cos\theta_{n}-\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}$ converges. Therefore,
$x_{n}=\beta_{n}\Big{[}\frac{\alpha_{n}}{\beta_{n}}\big{(}1-\cos\theta_{n}+\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\big{)}+\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n}\Big{]}$
converges implies that $\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n}\to 0$. Since $g$ is unipotent, it follows that $g=I$.
Also, for any $t\in\mathbb{R}$, we can choose $\zeta_{n}\to 0$ so that $\beta_{n}(\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2})\sin\zeta_{n}\to t$. Then $\beta_{n}\frac{s_{n}}{a_{n}^{2}}\sin\zeta_{n}\to 0$ and $\beta_{n}\sin\zeta_{n}$ converges (to 00 or $t$ depending on if $s_{n}^{2}+a_{n}^{2}\to\infty$ or not). Since $\beta_{n}\to\infty$, we also get that $(\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2})\sin\zeta_{n}\to 0$. Since $\zeta_{n}\to 0$ and $\beta_{n}\to\infty$, we get that $\mathfrak{s}_{n}\mathfrak{a}_{n}Z_{n}\mathfrak{a}_{n}^{-1}\mathfrak{s}_{n}^{-1}\to I$ for $Z_{n}=\begin{pmatrix}\cos\zeta_{n}&\sin\zeta_{n}\\ -\sin\zeta_{n}&\cos\zeta_{n}\end{pmatrix}\in K$. Since $\zeta_{n}\to 0$, $(1-\cos\zeta_{n})/\sin\zeta_{n}\to 0$, hence $\alpha_{n}(1-\cos\zeta_{n})=(\beta_{n}\sin\zeta_{n})(\frac{\alpha_{n}}{\beta_{n}})(\frac{1-\cos\zeta_{n}}{\sin\zeta_{n}})\to 0$ and $-\alpha_{n}{\sin\zeta_{n}\over a_{n}^{2}}+\beta_{n}\big{(}1-\cos\zeta_{n}-{s_{n}\over a_{n}^{2}}\sin\zeta_{n}\big{)}={-\alpha_{n}\over\beta_{n}}{\beta_{n}\sin\zeta_{n}\over a_{n}^{2}}+\beta_{n}\sin\zeta_{n}({1-\cos\zeta_{n}\over\sin\zeta_{n}})-\beta_{n}{s_{n}\over a_{n}^{2}}\sin\zeta_{n}\to 0$. Thus,
$(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})\ Z_{n}\ (\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}\to(I,(t,0)).$
Now, replacing $\Theta_{n}$ by $\Theta_{n}Z_{n}^{-1}$, we may assume that $x_{n}\to 0$. Since $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\to 0$ and ${s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2}\not\to 0$, we get that $\sin\theta_{n}\to 0$. Replacing $\theta_{n}$ by a translate of $2k\pi$, we may assume that $\theta_{n}\to 0$. This implies that ${1-\cos\theta_{n}\ \over\sin\theta_{n}}\to 0$. Since $x_{n}\to 0$, we get that $\beta_{n}({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\to 0$, hence $\alpha_{n}(1-\cos\theta_{n}+{s_{n}\over a_{n}^{2}}\sin\theta_{n})\to 0$. Therefore,
$|\alpha_{n}{-\sin\theta_{n}\over a_{n}^{2}}|\leq|{\alpha_{n}\over\beta_{n}}||\beta_{n}({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}|\to 0$
and
$\begin{array}[]{cl}&|\beta_{n}(1-\cos\theta_{n}-{s_{n}\over a_{n}^{2}}\sin\theta_{n})|\\ =&|{1-\cos\theta_{n}\over\sin\theta_{n}}{a_{n}^{2}\over s_{n}^{2}+a_{n}^{4}}-{s_{n}\over s_{n}^{2}+a_{n}^{4}}||\beta_{n}({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}|\to 0.\end{array}$
This implies that $H=V_{0}=\{(I,(x,0))\mid x\in\mathbb{R}\}$.
Case $2$: Suppose ${\beta_{n}\over\alpha_{n}}\to 0$. Let $F=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Since $(g_{n},v_{n})K(g_{n},v_{n})^{-1}\to H$, conjugating by $F$, we get that $(Fg_{n},Fv_{n})K(Fg_{n},Fv_{n})^{-1}\to FHF^{-1}$. Since $Fv$ satisfies case (i) we get that $FHF^{-1}=V_{0}$. Thus, $H=V_{\infty}=\{(I,(0,y)\mid y\in\mathbb{R}\}$.
Case $3$: Suppose $c\not=0$ is a limit point of $|{\alpha_{n}\over\beta_{n}}|$. By passing to a subsequence we may assume that ${\alpha_{n}\over\beta_{n}}\to c$. There are two cases here: if one of $a_{n}$ or $s_{n}$ diverges to $\infty$, then as in case (i) we can show that $H=V_{0}$. We now assume that $a_{n}=1$ and $s_{n}=0$. Then, choosing $\zeta_{n}\rightarrow 0$ such that $\beta_{n}\sin\zeta_{n}\rightarrow t$ for any preassigned real number $t$, it is easy to see that
$g_{n}Z_{n}g_{n}^{-1}\rightarrow\big{(}I,\ (t,-ct)\big{)}.$
for suitable $Z_{n}\in K$. Therefore, in this case as in case (i) we may show that $H=V_{c}=\{(I,(t,ct))\mid t\in\mathbb{R}\}$.
We now consider the case when $(v_{n})$ has a convergent subsequence. In this case, by passing to a subsequence and conjugating by $(I,-v_{n})$, we may assume that $(\mathfrak{s}_{n}\mathfrak{a}_{n},0)K(\mathfrak{s}_{n}\mathfrak{a}_{n},0)^{-1}\to H$. Since $\mathfrak{s}_{n},\mathfrak{a}_{n}\in SL(2,\mathbb{R})$ and $K\subset SL(2,\mathbb{R})$, $H\subset SL(2,\mathbb{R})$. Recall that we also have $s_{n}\to\infty$ or $s_{n}=0$ and $a_{n}\to\infty$ or $a_{n}=1$. If both $s_{n}=0$ and $a_{n}=1$, we are done. So we may assume that either $s_{n}\to\infty$ or $a_{n}\to\infty$. We consider the case that $s_{n}\to\infty$ and other case that $a_{n}\to\infty$ is similar. We have
$T_{n}=\begin{pmatrix}\cos\theta_{n}-{s_{n}\over a_{n}^{2}}\sin\theta_{n}&({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\\ {-\sin\theta_{n}/a_{n}^{2}}&\cos\theta_{n}+{s_{n}\over a_{n}^{2}}\sin\theta_{n}\end{pmatrix}\to g=\begin{pmatrix}g_{1,1}&g_{1,2}\\ g_{2,1}&g_{2,2}\end{pmatrix},{~{}~{}\rm say}.$
This in particular, implies that $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}$ converges. Since ${s_{n}\over a_{n}^{2}}+a_{n}^{2}\to\infty$, $\sin\theta_{n}\to 0$ and since ${\sin\theta_{n}\over a_{n}^{2}}\to g_{2,1}$, we get that $g_{2,1}=0$. Also, for $t\in\mathbb{R}$, we choose $\zeta_{n}\to 0$ such that $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\zeta_{n}\to t$. Then $\cos\zeta_{n}\to 1$ and ${s_{n}\over a_{n}^{2}}\sin\zeta_{n}\to 0$, hence $g_{n}Z_{n}g_{n}^{-1}\to\begin{pmatrix}1&t\\ 0&1\end{pmatrix}$ for suitable $Z_{n}\in K$. If, in addition $g_{1,2}=0$, then $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\to 0$ and hence ${s_{n}\over a_{n}^{2}}\sin\theta_{n}\to 0$. Thus, $\cos\theta_{n}$ converges to $g_{1,1}$ as well as $g_{2,2}$ but $g_{1,1}g_{2,2}=1$. Therefore, $g_{1,1}=\pm 1=g_{2,2}$. In fact, taking $\theta_{n}=\pi$, we get that
$g_{n}\Theta_{n}g_{n}^{-1}=\begin{pmatrix}\cos\theta_{n}-{s_{n}\over a_{n}^{2}}\ \sin\theta_{n}&({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\ \sin\theta_{n}\\ -\sin\theta_{n}/a_{n}^{2}&\cos\theta_{n}+{s_{n}\over a_{n}^{2}}\sin\theta_{n}\end{pmatrix}=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}.$
Thus, we have $H=\tilde{N^{+}}$, that is $g_{n}Kg_{n}^{-1}\rightarrow\tilde{N^{+}}$.
5.3. Limits of maximal diagonalizable subgroups
Now we study limit points of the set of maximal diagonalizable subgroups of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. We consider the conjugacy limits $(g_{n},v_{n})A(g_{n},v_{n})^{-1}$ of $A$ for different types of $g_{n}$ and $v_{n}$.
Proof of Theorem 1.9.
Let $A=\left\{\begin{pmatrix}a&0\\ 0&1/a\end{pmatrix}\mid a\in\mathbb{R}^{+}\right\}$ and $H$ be a limit point of $\{(g,v)A(g,v)^{-1}\mid g\in SL(2,\mathbb{R}),~{}v\in\mathbb{R}^{2}\}$. So, there are $g_{n}\in SL(2,\mathbb{R})$ and $v_{n}\in\mathbb{R}^{2}$ such that $(g_{n},v_{n})A(g_{n},v_{n})^{-1}\to H$. Using Iwasawa decomposition, we may assume that $g_{n}\in N^{+}$. Let $g_{n}=\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix}$ and $v_{n}=(\alpha_{n},\beta_{n})$. Suppose $(g,v)\in H$. Then, there are $a_{n}\in\mathbb{R}^{+}$ such that $(g_{n},v_{n})\mathfrak{a}_{n}(g_{n},v_{n})^{-1}\to(g,v)$, where $\mathfrak{a}_{n}=\begin{pmatrix}a_{n}&0\\ 0&1/a_{n}\end{pmatrix}$. This implies that
$\begin{pmatrix}a_{n}&({1\over a_{n}}-a_{n})s_{n}\\ 0&{1\over a_{n}}\end{pmatrix}\to g,~{}~{}{\rm and}~{}~{}$
$((1-a_{n})\alpha_{n}+(a_{n}-{1\over a_{n}})s_{n}\beta_{n},(1-{1\over a_{n}})\beta_{n})\to v.$
By passing to subsequences and taking conjugates, we may assume that $s_{n}=0$ or $s_{n}\to\infty$, $\alpha_{n}=0$ or $\alpha_{n}\to\infty$ and $\beta_{n}=0$ or $\beta_{n}\to\infty$.
We first consider the case that $s_{n}\to\infty$. If both $\alpha_{n}=0$ and $\beta_{n}=0$, then it is easy to see that $H=N^{+}=\left\{\begin{pmatrix}1&t\\ 0&1\end{pmatrix}\mid t\in\mathbb{R}\right\}$. So assume that either $\alpha_{n}\to\infty$ or $\beta_{n}\to\infty$. Then $v_{n}\to\infty$. This implies by Lemma 5.1 that $g$ is unipotent. Therefore, $a_{n}\to 1$.
Since $g_{n}\in N^{+}$, and $B=AN^{+}$, it follows that $(g_{n},v_{n})A(g_{n},v_{n})^{-1}$ are closed subgroups of $B\ltimes\mathbb{R}^{2}$. Since $B\ltimes\mathbb{R}^{2}$ is isomorphic to a group of upper triangular matrices with positive diagonal entries, $B\ltimes\mathbb{R}^{2}$ is an exponential Lie group. Hence, by Proposition 2.2, $H$ is a connected one-dimensional subgroup of $N^{+}\ltimes\mathbb{R}^{2}$ which is the $3$-dimensional Heisenberg group. Therefore,
$H=V_{(a,b,c)}=\left\{\left(\begin{pmatrix}1&at\\ 0&1\end{pmatrix},\left(ct+{abt^{2}\over 2},bt\right)\right)\mid t\in\mathbb{R}\right\}$
for some $a,b,c\in\mathbb{R}$. Then for each $t\in\mathbb{R}$, there are $b_{n}$ (depending on $t$) such that
$b_{n}\to 1,~{}~{}({1\over b_{n}}-b_{n})s_{n}\to at~{}~{}{\rm and}~{}~{}(1-{1\over b_{n}})\beta_{n}\to bt.$
Case $1$: Suppose ${s_{n}\over\beta_{n}}\to r\in\mathbb{R}$. Then $a=-2br$ and hence
$(1-b_{n})\alpha_{n}-{(1+b_{n})\over b_{n}}(1-b_{n})s_{n}\beta_{n}\to ct-rb^{2}t^{2}.$
If $b=0$, then $H=V_{0}$. So, we may assume that $b\not=0$. Also
$2(1-b_{n})s_{n}\beta_{n}-{(1+b_{n})\over b_{n}}(1-b_{n})s_{n}\beta_{n}\to-rb^{2}t^{2}.$
Therefore,
$(1-b_{n})\alpha_{n}-2(1-b_{n})s_{n}\beta_{n}\to ct.$
This implies that $2s_{n}-{\alpha_{n}\over\beta_{n}}\to{c\over b}$. Let $d=\lim 2s_{n}-{\alpha_{n}\over\beta_{n}}$. Then $c=db$. Thus, $H=V_{(r_{1}b,r_{2}b,b)}$ for some $r_{1},r_{2}\in\mathbb{R}$.
Case $2$: Suppose ${\beta_{n}\over s_{n}}\to p\in\mathbb{R}$, Then $b=-{pa\over 2}$ and hence
$(1-b_{n})\alpha_{n}-{(1+b_{n})\over b_{n}}(1-b_{n})s_{n}\beta_{n}\to ct-{pa^{2}t^{2}\over 4}.$
If $a=0$, then $H=V_{0}$. So, we may assume that $a\not=0$. Thus,
$(1-b_{n})\alpha_{n}-2(1-b_{n})s_{n}\beta_{n}\to ct.$
This implies that ${\alpha_{n}\over s_{n}}-2\beta_{n}\to 2{c\over a}$, hence $c={da\over 2}$ where $d=\lim{\alpha_{n}\over s_{n}}-2\beta_{n}$. Thus, $H=V_{(a,r_{1}a,r_{2}a)}$ for some $r_{1},r_{2}\in\mathbb{R}$.
We now consider the case $s_{n}=0$. Then $g=I$ and
$((1-a_{n})\alpha_{n},(1-{1\over a_{n}})\beta_{n})\to v$
for some $a_{n}\to 1$.
Case $3$: Suppose ${\beta_{n}\over\alpha_{n}}\to 0$. Then since $v_{n}=(\alpha_{n},\beta_{n})\to\infty$, $\alpha_{n}\to\infty$ and $(1-{1\over a_{n}})\beta_{n}={-1\over a_{n}}(1-a_{n})\alpha_{n}{\beta_{n}\over\alpha_{n}}\to 0$. Therefore $H\subset V_{0}$. For any $t\in\mathbb{R}$, taking $b_{n}=1-{t\over\alpha_{n}}$, we get that $b_{n}\to 1$, $(1-b_{n})\alpha_{n}=t$. Therefore, $((1-b_{n})\alpha_{n},(1-{1\over b_{n}})\beta_{n})\to(t,0)$. Thus, $H=V_{0}$.
Case $4$: Suppose ${\alpha_{n}\over\beta_{n}}\to 0$. Then $(1-a_{n})\alpha_{n}=-a_{n}(1-{1\over a_{n}})\beta_{n}{\alpha_{n}\over\beta_{n}}\to 0$. Therefore, $H\subset V_{\infty}=\{(0,t)\mid t\in\mathbb{R}\}$. For any $t\in\mathbb{R}$, take $b_{n}={\beta_{n}\over\beta_{n}-t}$. Then $b_{n}\to 1$ and $((1-b_{n})\alpha_{n},(1-{1\over b_{n}})\beta_{n})\to(0,t)$. Thus, $H=V_{\infty}$.
Case $5$: Suppose ${\beta_{n}\over\alpha_{n}}\to c\not=0$. Let $x,y\in\mathbb{R}$ be such that $v=(x,y)$. Then $(1-a_{n})\alpha_{n}\to x$ and $(1-{1\over a_{n}})\beta_{n}\to y$. Since $(1-{1\over a_{n}})\beta_{n}={-1\over a_{n}}(1-a_{n})\alpha_{n}{\beta_{n}\over\alpha_{n}}$, we get that $y=-cx$. Therefore $H\subset V_{r}=\{(t,rt)\mid t\in\mathbb{R}\}$ for $r=-c$.
For $t\in\mathbb{R}$, let $b_{n}=1-{t\over\alpha_{n}}$. Then $b_{n}\to 1$ and $(1-b_{n})\alpha_{n}=t$. Also, $(1-{1\over b_{n}})\beta_{n}={-t\over\alpha_{n}-t}\beta_{n}\to-tc$. Therefore $V_{r}\subset H$. Thus, $H=V_{r}$.
5.4. Limits of unipotent and Borel subgroups
We now find the conjugacy limits of the Borel subgroup
$B:=\left\{\left(\begin{pmatrix}a&s\\ 0&1/a\end{pmatrix}:\ a,s\in\mathbb{R}\right)\right\}$
of $\operatorname{SL}(2,\mathbb{R})$ inside $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
Proof of Theorem 1.10.
Clearly, a conjugacy limit of $B$, when conjugated by elements of the type $(g,0)$ with $g\in\operatorname{SL}(2,\mathbb{R})$, is $B$ itself or a conjugate of $B$ by elements of $K$. For conjugation by more general type of elements, it is enough to consider conjugation by elements of the form $(I,v)$ with $v\in\mathbb{R}^{2}$. So, let $\mathfrak{g}_{n}=(I,v_{n})$ and $\mathfrak{b}_{n}=\begin{pmatrix}a_{n}&s_{n}\\ 0&1/a_{n}\end{pmatrix}$. Then a simple calculation shows that
$\mathfrak{g}_{n}\mathfrak{b}_{n}\mathfrak{g}_{n}^{-1}=\left(\begin{pmatrix}a_{n}&s_{n}\\ 0&1/a_{n}\end{pmatrix},\Big{(}(1-a_{n})\alpha_{n}-s_{n}\beta_{n},\ (1-1/a_{n})\beta_{n}\Big{)}\right)$
$=\big{(}\mathfrak{b}_{n},(x_{n},y_{n})\big{)},\ (\text{say})$
where $v_{n}=(\alpha_{n},\beta_{n})$. Assume that $v_{n}\to\infty$ and $\mathfrak{g}_{n}\mathfrak{b}_{n}\mathfrak{g}_{n}^{-1}\to(g,v)$. Then by Lemma 5.1, $g$ is unipotent, hence $a_{n}\rightarrow 1$ necessarily. We shall consider the following cases.
Case $1$: Suppose $\frac{\alpha_{n}}{\beta_{n}}\rightarrow 0$. Then $\beta_{n}\to\infty$. If $s_{n}\rightarrow s$ for some non-zero $s$, then it is easy to see that $x_{n}$ cannot converge. Hence, $s_{n}\rightarrow 0$. Given $x,y\in\mathbb{R}$, choose $s_{n}\rightarrow 0$ and $a_{n}\rightarrow 1$ in such a way that $-\beta_{n}s_{n}\rightarrow x$ and $(1-1/a_{n})\beta_{n}\rightarrow y$. Then $(1-a_{n})\alpha_{n}=-a_{n}(1-{1\over a_{n}})\beta_{n}{\alpha_{n}\over\beta_{n}}\to 0$, hence $\mathfrak{g}_{n}\mathfrak{b}_{n}^{\prime}\mathfrak{g}_{n}^{-1}\rightarrow\big{(}I,(x,y)\big{)}$ for suitable $\mathfrak{b}_{n}^{\prime}\in B$. This shows that, in this case, the limit subgroup is $\mathbb{R}^{2}$.
Case $2$: Suppose $\frac{\beta_{n}}{\alpha_{n}}\rightarrow 0$. Then $\alpha_{n}\to\infty$. Let $x\in\mathbb{R}$. Then for $\tilde{\mathfrak{b}}_{n}=\begin{pmatrix}1-{x\over\alpha_{n}}&0\\ 0&{\alpha_{n}\over\alpha_{n}-x}\end{pmatrix}$, we get that $\mathfrak{g}_{n}\tilde{\mathfrak{b}_{n}}\mathfrak{g}_{n}^{-1}\to(I,(t,0))$. Thus, $\{(I,(t,0))\mid t\in\mathbb{R}\}\subset H$.
If $\beta_{n}$ converges, then $y_{n}=(1-{1\over a_{n}})\beta_{n}\to 0$ and it can be shown that $H=N^{+}\ltimes\{(t,0)\mid t\in\mathbb{R}\}$. So, we may assume that $\beta_{n}\to\infty$. Since $\mathfrak{b}_{n}$ converges, $s_{n}$ converges, say to $s$. Since $\beta_{n}\to\infty$ and $x_{n}$ converges, we get that $(1-a_{n}){\alpha_{n}\over\beta_{n}}\to s$.
If ${\beta_{n}^{2}\over\alpha_{n}}\to c\in\mathbb{R}$, then $y_{n}={-1\over a_{n}}(1-a_{n}){\alpha_{n}\over\beta_{n}}{\beta_{n}^{2}\over\alpha_{n}}\to-sc$. Therefore, by Proposition 2.2 we get that $H=\{(\mathfrak{s},(t,-cs)\mid t,s\in\mathbb{R}\}$ where $\mathfrak{s}=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}$.
If ${\beta_{n}^{2}\over\alpha_{n}}\to\infty$, then $(1-a_{n}){\alpha_{n}\over\beta_{n}}=(1-{1\over a_{n}})\beta_{n}(-a_{n}){\alpha_{n}\over\beta_{n}^{2}}\to 0$. Therefore, by Proposition 2.2 we get that $H=\mathbb{R}^{2}=\{(I,(t,r)\mid t,r\in\mathbb{R}\}$.
Case $3$: Suppose $\frac{\alpha_{n}}{\beta_{n}}\rightarrow c$ for some non-zero real number $c$. Note that we are necessarily assuming both $|\alpha_{n}|$ and $|\beta_{n}|$ are unbounded. Otherwise, the limit subgroup will be a conjugate of $B$. Now, suppose $x_{n}\rightarrow x$. Then $\frac{x_{n}}{\beta_{n}}\rightarrow 0$. But,
$\frac{x_{n}}{\beta_{n}}=\frac{(1-a_{n})\alpha_{n}}{\beta_{n}}-s_{n}$
which implies that $s_{n}\rightarrow 0$ in this case as well. Therefore $H\subset\mathbb{R}^{2}$. Thus, by Proposition 2.2, $H$ is of dimension $2$ and hence $H=\mathbb{R}^{2}$.
We next determine the conjugacy limits of the unipotent subgroup $N^{+}$.
Proof.
Proof of Theorem 1.11 Since $N^{+}$ is normal in $B$, using Iwasawa decomposition in $SL(2,\mathbb{R})$, it follows that whenever $g_{n}\in\operatorname{SL}(2,\mathbb{R})$, the limit of $g_{n}N^{+}g_{n}^{-1}$ is either $N^{+}$ or its conjugates by elements of $K$. For conjugation by other elements of $G$, without loss of generality, we may consider conjugation by elements of the form $\mathfrak{g}_{n}=(k_{n},v_{n})$ with $k_{n}\in K$ and $v_{n}=(\alpha_{n},\beta_{n})\in\mathbb{R}^{2}$. Also, it is easy to see that if $L$ is a closed subgroup of $G$, $g_{n}Lg_{n}^{-1}\rightarrow H$ and $k_{n}\rightarrow k$, then $k_{n}g_{n}Lg_{n}^{-1}k_{n}^{-1}\rightarrow kHk^{-1}$. In view of this, it is enough to consider conjugation of $N^{+}$ by elements of the form $\mathfrak{g}_{n}=(I,v_{n})$.
Now, suppose $\mathfrak{g}_{n}=\big{(}I,v_{n}\big{)}$ with $v_{n}=(\alpha_{n},\beta_{n})$. Then for any $\mathfrak{s}_{n}=\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix}$,
$\mathfrak{g}_{n}\mathfrak{s}_{n}\mathfrak{g}_{n}^{-1}=\left(\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix},\big{(}-s_{n}\beta_{n},0\big{)}\right).$
If $|\beta_{n}|\rightarrow\infty$, then choosing $s_{n}\rightarrow 0$ suitably $\big{(}I,(t,0)\big{)}$ can be realized as a limit of $\mathfrak{g}_{n}\mathfrak{s}_{n}\mathfrak{g}_{n}^{-1}$ for any real number $t$. Also, if $\mathfrak{g}_{n}\mathfrak{s}_{n}\mathfrak{g}_{n}^{-1}$ converges, then $s_{n}$ as well as $s_{n}\beta_{n}$ converges. but, as $|\beta_{n}|\to\infty$, we have $s_{n}\to 0$. This shows that, in this case, the limit of $\mathfrak{g}_{n}N^{+}\mathfrak{g}_{n}^{-1}$ is $V_{0}$.
Acknowledgement .
The first author thanks the Indian Satitistical Institute Bangalore Center for its hospitality during his several visits while this work was going on. The second author acknowledges the support from SERC through the grant under MATRICS MTR/2022/000429.

### v2

@@
-IRS on orbits and the case of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$
+Invariant random subgroups on certain orbits
 Abstract.
@@
 In this article, we investigate the existence of invariant random subgroups of $G$ supported on various orbits of the conjugation action of $G$ on $\text{Sub}_{G}$.
-We also find the orbit closure for each of these orbits in the case of the Lie group $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-Key words and phrases:
-Connected Lie groups, spaces of closed subgroups, Chabauty topology, invariant random subgroups, conjugacy limits
-1991 Mathematics Subject Classification:
-Primary: $22$E15151515, $28$D15151515; secondary: $37$A15151515, $37$B02020202
 1. Introduction
-Let $G$ be a connected (real) Lie group and ${\rm Sub}_{G}$ be the space of (closed) subgroups of $G$ equipped with the Chabauty topology introduced by Claude Chabauty in [4].
-The Chabauty topology on SubG is generated by the subsets $\{H\in{\rm Sub}_{G}\mid H\cap K=\emptyset\}$, $\{H\in{\rm Sub}_{G}\mid H\cap U\not=\emptyset\}$, where $K\subset G$ is compact and $U\subset G$ is open.
+Let $G$ be a connected (real) Lie group and $\text{Sub}_{G}$ be the space of (closed) subgroups of $G$ equipped with the Chabauty topology introduced by Claude Chabauty in [6].
+The Chabauty topology on $\text{Sub}_{G}$ is generated by the subsets $\{H\in{\rm Sub}_{G}\mid H\cap K=\emptyset\}$, $\{H\in{\rm Sub}_{G}\mid H\cap U\not$$\emptyset\}$, where $K\subset G$ is compact and $U\subset G$ is open.
 With respect to the Chabauty topology, $\text{Sub}_{G}$ is a compact metrizable space (cf.
 Lemma E.1.1 of [4]).
-Moreover, a sequence $H_{n}\in\text{Sub}_{G}$ converges to $H\in\text{Sub}_{G}$ if and only if for any $h\in H$ there is a sequence $h_{n}\in H_{n}$ such that $h_{n}$ converges to $h$ and for any sequence $x_{k_{n}}\in H_{k_{n}}$, with $k_{n+1}>k_{n}$ and $x_{k_{n}}\to x$, we have $x\in H$ (cf.
+Moreover, a sequence $H_{n}\in\text{Sub}_{G}$ converges to $H\in\text{Sub}_{G}$ if and only if for any $h\in H$ there is a sequence $h_{n}\in H_{n}$ such that $h_{n}$ converges to $h$, and for any sequence $x_{k_{n}}\in H_{k_{n}}$, with $k_{n+1}>k_{n}$ and $x_{k_{n}}\to x$, we have $x\in H$ (cf.
 Lemma E.1.2 of [4]).
-The group $G$ acts on SubG by conjugation: $(g,H)\mapsto gHg^{-1}$.
-An invariant random subgroup (hereafter IRS) is a probability measure $\mu$ on SubG that is invariant under the conjugation action of $G$ on SubG.
-We say that a Borel subset ${\mathcal{O}}$ of SubG supports an IRS $\mu$ if $\mu({{\mathcal{O}}})=1$: a Borel subset ${{\mathcal{O}}}$ supporting an IRS $\mu$ need not be the support of $\mu$ but $\overline{{\mathcal{O}}}$ contains the support of $\mu$.
-Dirac measure supported on a closed normal subgroup of a connected Lie group is an example of an IRS.
-Suppose $\Gamma$ is a lattice inside the connected Lie group $G$, i.e., $\Gamma$ is a discrete subgroup of $G$ such that $G/\Gamma$ carries a probability measure $\mu$ that is invariant under the natural action of $G$ on $G/\Gamma$.
-Then the map from $G$ to $\text{Sub}_{G}$ sending $g$ to $g\Gamma g^{-1}$ factors through $G/\Gamma$, and the push forward of $\mu$ by this map becomes a conjugation-invariant probability measure supported on conjugacy classes of $\Gamma$.
-So, invariant random subgroups can be thought of as a generalization of both normal subgroups and lattices.
-Invariant random subgroups arise naturally as stabilizers of measure preserving actions of $G$ on a probability space $(X,\mu)$.
-By a result of Varadarajan ([14]), the stabilizer of almost every point of this action is closed in $G$.
-Moreover, the stabilizer map from $X$ to $\text{Sub}_{G}$ given by $x\to G_{x}$ is measurable.
-Then, one can push the measure $\mu$ forward to an IRS of $G$.
-In fact, every IRS can be obtained this way (see [1], [2] for more details).
-In recent times, IRS has attracted many studies (see for examples [1] and [2]).
+The group $G$ acts on $\text{Sub}_{G}$ by conjugation: $(g,H)\mapsto gHg^{-1}$.
+An invariant random subgroup (hereafter IRS) is a probability measure $\mu$ on $\text{Sub}_{G}$ that is invariant under the conjugation action of $G$ on $\text{Sub}_{G}$.
+Dirac measure supported on a closed normal subgroup is an IRS.
+Also, if $\Gamma$ is a lattice in $G$, i.e., $\Gamma$ is a discrete subgroup of $G$ such that $G/\Gamma$ carries a $G$-invariant probability measure $\mu$, then the push forward of $\mu$ under the map $g$ to $g\Gamma g^{-1}$ defines an IRS on $G$.
+Thus, invariant random subgroups can be thought of as a generalization of both normal subgroups and lattices.
+The study of invariant random subgroups and ergodic theory are deeply intertwined as an IRS can be realized as stabilizers of a measure-preserving action of the group (see [1] and [3] for more details).
+We say that a Borel subset $\mathcal{O}$} of $\text{Sub}_{G}$ supports an IRS $\mu$ if $\mu({{\mathcal{O}}})=1$: a Borel subset $\mathcal{O}$}} supporting an IRS $\mu$ need not be the support of $\mu$ but $\overline{{\mathcal{O}}}$ contains the support of $\mu$.
+In recent times, IRS has attracted many studies (cf.
+[1], [2], [3], [5], [9] and the references therein).
 In this article, we are interested in the existence of invariant random subgroups supported on certain orbits of the conjugation action of $G$ on $\text{Sub}_{G}$.
-We start with the orbit of a semisimple Levi subgroup.
-Let $G$ be a connected Lie group with a closed semisimple Levi subgroup $L$; it may be noted that semisimple Levi subgroups of connected Lie groups need not be closed.
-But, Levi subgroups of linear groups are closed.
+We first consider the orbit of semisimple Levi subgroups.
+Let $G$ be a connected Lie group with a closed semisimple Levi subgroup $L$; it may be noted that semisimple Levi subgroups of connected Lie groups need not be closed, but, Levi subgroups of linear groups are closed.
 Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$.
@@
     the semisimple Levi subgroup $L$ is normal in $G$.
-In particular, the set of semisimple Levi subgroups $L_{G}$ is closed in SubG if and only if $L$ is normal.
+In particular, the set of semisimple Levi subgroups $L_{G}$ is closed in $\text{Sub}_{G}$ if and only if $L$ is normal.
 Corollary 1.2.
 Any Levi IRS of a connected Lie group $G$ is a Dirac measure supported on a Levi subgroup of $G$.
-So, any connected Lie group $G$ admitting a Levi IRS is of the form $G=L\times S$ where $L$ is semisimple and $S$ is solvable.
-In particular, $G=\operatorname{SL}(2,\mathbb{R})\times\mathbb{R}^{2}$ admits a Levi IRS, whereas $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ does not admit a Levi IRS.
+Thus, any connected Lie group $G$ admitting a Levi IRS is of the form $G=L\times S$ where $L$ is semisimple and $S$ is solvable.
+In particular, for $n\geq 2$, $\operatorname{SL}(n,\mathbb{R})\times\mathbb{R}^{n}$ admits a Levi IRS, whereas $\operatorname{SL}(n,\mathbb{R})\ltimes\mathbb{R}^{n}$ does not admit a Levi IRS.
 As in the case of Levi subgroups, the set of all maximal compact subgroups of a connected Lie group is also the orbit of a maximal compact subgroup under the conjugation action.
@@
 Let $G$ be a connected real algebraic group and $B_{G}$ be the set of all Borel subgroups of $G$.
-Then $B_{G}$ is closed in SubG and there is an IRS supported on $B_{G}$ if and only if $G$ is amenable.
+Then $B_{G}$ is closed in $\text{Sub}_{G}$, and there is an IRS supported on $B_{G}$ if and only if $G$ is amenable.
 Example 1.5.
@@
 Then $B=SO(2,\mathbb{R})\ltimes R^{3}$ is a Borel subgroup of $G$ and the normalizer of $B$ is $O(2,\mathbb{R})\ltimes\mathbb{R}^{3}$.
-Thus, $SO(3,\mathbb{R})$-invariant measure on the quotient space $SO(3,\mathbb{R})/O(2,\mathbb{R})$ induces an IRS on the set of Borel subgroups $B_{G}$ of $G$.
+Thus, any $SO(3,\mathbb{R})$-invariant measure on the quotient space $SO(3,\mathbb{R})/O(2,\mathbb{R})$ induces an IRS on the set of Borel subgroups $B_{G}$ of $G$.
 We next consider the orbit of a maximal diagonalizable subgroup $D$.
@@
 Theorem 1.6.
-Let $G$ be a connected real algebraic group and $D_{G}$ be the set of all maximal diagnolizable subgroups of $G$. Then there is an IRS supported on $D_{G}$ if and only if $D$ is central in $G$.
-We have seen that the group $\operatorname{SL}(2,\mathbb{R})\times\mathbb{R}^{2}$ admits a Levi IRS, whereas, the group $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ does not admit a Levi IRS.
-In particular, the orbit of a semisimple Levi subgroup is not closed in $\text{Sub}_{G}$ when $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-This leads to a natural question of studying the closure of the orbit under the conjugation action of a semisimple Levi subgroup.
-The same question can be asked for other orbits discussed above as well.
-In this article, we address these questions in the case of the Lie group $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-Note that, studying these questions is the same as investigating about conjugacy limits of various subgroups of $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-The study of conjugacy limits is an interesting problem in its own right and has been an active area of research aimed at understanding the space $\text{Sub}_{G}$.
-For example, in [10], [11], the author has studied the conjugacy limits of diagonal subgroups in $\operatorname{SL}(3,\mathbb{R})$ and $\operatorname{SL}(n,\mathbb{R})$ respectively.
-One may also look at the references cited in these two papers for related works and their relation to geometry.
-Given a closed subgroup $H$ of $G$, by a conjugacy limit of $H$ in $\text{Sub}_{G}$, we mean the limit of $(g_{n}Hg_{n}^{-1})$ for a sequence $(g_{n})$ in $G$.
-Note that $(g_{n}Hg_{n}^{-1})$ always has a convergent subsequence as $\text{Sub}_{G}$ is compact.
-The following theorem gives a complete description of all the limit points of the set $L_{G}$.
-We write any element of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ as $(g,v)$ with $g\in\operatorname{SL}(2,\mathbb{R})$ and $v\in\mathbb{R}^{2}$.
-Also, for any two elements $(g,v)$ and $(g^{\prime},v^{\prime})$, the group operation is defined as follows:
-$(g,v)(g^{\prime},v^{\prime})=(gg^{\prime},v+g(v^{\prime})).$
-We identify $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ as closed subgroups of $G$ by the maps $g\mapsto(g,(0,0))$ and $v\mapsto(I,v)$ where $I$ denotes the $2\times 2$ identity matrix.
-For any $c\in\mathbb{R}$, we define a unipotent subgroup $N^{+}_{c}$ by
-$N^{+}_{c}=\left\{\big{(}\begin{pmatrix}1&x\\ 0&1\end{pmatrix},(cx,0)\big{)}:x\in\mathbb{R}\right\}.$
-For $c=0$, we denote $N^{+}_{c}$ by $N^{+}$, which is identified (and isomorphic) to the group of unipotent upper triangular matrices in $\operatorname{SL}(2,\mathbb{R})$.
-Theorem 1.7.
-Any subgroup in $\overline{L_{G}}$ is either a Levi subgroup of $G$ or a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
-Next, we consider the conjugacy limit of maximal compact subgroups, that is, the closure of $K_{G}=\{gKg^{-1}\mid g\in G\}$ for any maximal compact subgroup $K$ of $G=SL(2,\mathbb{R})\times\mathbb{R}^{2}$.
-For example, we can take $K=SO(2,\mathbb{R})$.
-For this purpose, we introduce the following subgroups of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$:
-$V_{c}=\{(I,(t,ct))\mid t\in\mathbb{R}\},\ c\in\mathbb{R},~{}~{}V_{\infty}=\{(I,(0,t))\mid t\in\mathbb{R}\}$
-and
-$\tilde{N}^{+}=\left\{\pm\begin{pmatrix}1&s\\ 0&1\end{pmatrix}:s\in\mathbb{R}\right\}.$
-It may be noted that $V_{c},c\in\mathbb{R}\cup\{\infty\}$ is the set of all one-dimensional subspaces of $\mathbb{R}^{2}$, the projective space ${\mathbb{P}}^{1}$.
-Theorem 1.8.
-The closure of $K_{G}$ consists of conjugates of the subgroups $V_{c},c\in\mathbb{R}\cup\{\infty\}$ and $\tilde{N^{+}}$, in addition to the conjugates of $K$.
-The next result identifies limit points of the set of maximal diagonalizable subgroups of $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-Let $A=\left\{\begin{pmatrix}a&0\\ 0&\frac{1}{a}\end{pmatrix}\mid a\in\mathbb{R}^{+}\right\}$ and $D_{G}=\{gAg^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}A$.
-We now introduce the following one-dimensional subgroups
-$V_{(a,b,c)}=\left\{\left(\begin{pmatrix}1&at\\ 0&1\end{pmatrix},\left(ct+\frac{abt^{2}}{2},bt\right)\right)\mid t\in\mathbb{R}\right\}$
-for any $a,b,c\in\mathbb{R}$.
-It may be noted that $V_{a,b,c}$’s are the one-parameter subgroups in the $3$-dimensional Heisenberg group and $V_{(1,0,c)}=N^{+}_{c}$, $V_{(0,1,c)}=V_{c}$ and $V_{(0,0,1)}=V_{\infty}$.
-Theorem 1.9.
-The closure of $D_{G}$ consists of conjugates of $V_{(r_{1},r_{2},r_{3})}$ for any $r_{1},r_{2},r_{3}\in\mathbb{R}$ with $r_{1}=1$ or $r_{3}=1$ or $r_{1}=0$ and $r_{2}=1$ in addition to the conjugates of the diagonal group $A$.
-Let $B$ denote the group of all upper triangular matrices in $\operatorname{SL}(2,\mathbb{R})$.
-Then $B$ is a Borel subgroup of $\operatorname{SL}(2,\mathbb{R})$ and $B\ltimes\mathbb{R}^{2}$ is a Borel subgroup of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-Then, using the Iwasawa decomposition in $\operatorname{SL}(2,\mathbb{R})$, it is easy to see that, for conjugacy limits of $B\ltimes\mathbb{R}^{2}$, we only need to consider conjugation by elements of $K=\operatorname{SO}(2,\mathbb{R})$, hence conjugates of $B$ form a closed subset (but by Theorem 1.4 no IRS is supported on the conjugates of $B$ as $G$ is not amenable).
-We therefore consider conjugacy limits of $B$ inside $G$ and identify all possible conjugacy limits of $B$
-Theorem 1.10.
-The closure of $\{gBg^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}$ consists of conjugates of $B$, $N^{+}\ltimes\{(t,0)\mid t\in\mathbb{R}\}~{}(\simeq\mathbb{R}^{2})$ and $\mathbb{R}^{2}$.
-The next result is concerned about the conjugacy limits of the unipotent subgroup $N^{+}$.
-Theorem 1.11.
-The closure of $\{gN^{+}g^{-1}\mid g\in SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}\}$ consists of conjugates of $N^{+}$ and $V_{0}$.
+Let $G$ be a connected real algebraic group and $D_{G}$ be the set of all maximal diagnolizable subgroups of $G$. Then the following are equivalent:
+  1. (1)
+    there is an IRS supported on $D_{G}$;
+  2. (2)
+    $D$ is central in $G$;
+  3. (3)
+    $D$ is normal in $G$.
+Moreover, if (1) holds, then any Levi subgroup of $G$ is compact and the solvable radical is a compact extension of its nilradical.
 2. Preliminary
-In this section, we prove certain technical results which are needed to discuss the existence or non-existence of IRS on orbits of the conjugation action of $G$ on $\text{Sub}_{G}$.
-Given a finite-dimensional vector space $V$, let ${\rm Gr}_{k}(V)$ be the be the Grassnannian manifold of $k$-dimensional subspaces of $V$.
-Equip $V$ with an inner product, and for a subspace $W$ of $V$, let $P_{W}$ denote the orthogonal projection onto the subspace $W$.
-We define a metric on ${\rm Gr}_{k}(V)$ using the operator norm ||.||||.||| | .
-| | of $P_{W}$’s as follows: for $W,W^{\prime}\in{\rm Gr}_{k}(V)$,
-$d(W,W^{\prime})=||P_{W}-P_{W^{\prime}}||.$
-The following Lemma is a comparison between the Grassmannian topology and the Chabauty topology of ${\rm Gr}_{k}(V)$ seen as a subset of $\text{Sub}_{V}$.
+In this section, we now collect a few results that connect the dynamics of $G$ on $\text{Sub}_{G}$ with projective linear actions on Grassmannians.
+Given a finite-dimensional vector space $V$, let ${\rm Gr}_{k}(V)$ be the Grassnannian manifold of $k$-dimensional subspaces of $V$.
+We include a proof of the following lemma because we are unable to locate it in the following form that compares the Grassmannian topology and the Chabauty topology of ${\rm Gr}_{k}(V)$ seen as a subset of $\text{Sub}_{V}$.
+Recall that the metric on ${\rm Gr}_{k}(V)$ is given by
+[…]
+for $W,W^{\prime}\in{\rm Gr}_{k}(V)$, with $P_{W}$ and $P_{W}^{\prime}$ being the orthogonal projections onto the subspaces $W$ and $W^{\prime}$ respectively.
+One may look at [9] for the measurability of the map.
 Lemma 2.1.
@@
 Then ${\rm Gr}_{k}(V)$ is homeomorphic to its image.
-In particular, the set of all subspaces of dimension $k$ is a closed set in Chaubauty topology.
+In particular, the set of all subspaces of dimension $k$ is a closed set in the Chaubauty topology.
 Proof.
 Suppose $W_{n}\to W$ in ${\rm Gr}_{k}(V)$.
-This implies that $||P_{W_{n}}-P_{W}||\rightarrow 0$ as $n\rightarrow\infty$.
+This implies that $P_{W_{n}}-P_{W}||\rightarrow 0$ as $n\rightarrow\infty$.
 Now let $w\in W$, we want to show that there exist $w_{n}\in W_{n}$ such that $w_{n}\rightarrow w$.
@@
 Then clearly $w_{n}\in W_{n}$ and
-$||w_{n}-w||=||P_{W_{n}}(w)-P_{W}(w)||\leq||P_{W_{n}}-P_{W}||||w||\rightarrow 0$
+[…]
 as $n\rightarrow\infty$.
@@
 Thus, $W_{n}\to W$ in SubV.
-This proves that the map $\pi$ is continuous.
+This proves that the map $\mathfrak{i}$ is continuous.
 Since ${\rm Gr}_{k}(V)$ is compact, ${\rm Gr}_{k}(V)$ is homeomorphic to $\mathfrak{i}({\rm Gr}_{k}(V))$.
 ∎
-It is known that if a sequence of closed subgroups $H_{n}$ of a Lie group $G$ converges to a closed subgroup $H$, then dim$(H)\geq{\rm limsup}~{}\text{dim}(H_{n})$ (cf.
-Theorem 1 of [15]).
-However, in the case of closed subgroups of thev same dimension and in the case of exponential Lie groups, we have the following.
+[…]
 Proposition 2.2.
-Let $G$ be a connected Lie group and $H_{n}\to H$ in SubG. Let $\mathcal{G}$, $\mathcal{H}_{n}$ and $\mathcal{H}$ be Lie algebras of $G$, $H_{n}$ and $H$ respectively.
-  1. (1)
+Let $G$ be a connected Lie group and $H_{n}\to H$ in $\text{Sub}_{G}$.
+Let $\mathcal{G}$, $\mathcal{H}_{n}$ and $\mathcal{H}$ be Lie algebras of $G$, $H_{n}$ and $H$ respectively.
 If $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$, then $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
-  2. (2)
-    If $G$ is an exponential Lie group and $H_{n}$’s are connected, then $H$ is also connected, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG and ${\rm limsup}~{}\text{dim}(H_{n})=\text{dim}(H)$.
 Remark 2.3.
-  1. (1)
 The map sending a closed subgroup to its Lie algebra is known to be a measurable map (cf.
 proof of Lemma 6 in [15]) but here we prove continuity under some restrictions though measurability is sufficient for our purpose.
-  2. (2)
-    It may be noted that $H$ need not always be connected even when $H_{n}$ are all connected and have the same dimension: e.g., cases $1$ and $2$ of Section $5.3$ provide a sequence of conjugate subgroups whose limit point has two connected components.
 Proof.
@@
 By passing to a subsequence, we may assume that $\mathcal{H}_{n}\to V$ in SubG.
-Then $V$ is a vector subspace.
-In is easy to see that it is a Lie subalgebra of $\mathcal{G}$ as well.
+Then $V$ is a vector subspace, and it is easy to see that $V$ is a Lie subalgebra of $\mathcal{G}$ as well.
 For $v\in V$, there are $v_{n}\in\mathcal{H}_{n}$ such that $v_{n}\to v$, hence $tv_{n}\to tv$ for all $t\in\mathbb{R}$.
@@
 Thus, $V\subset\mathcal{H}$.
-Assume that $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$.
-Then $\text{dim}(\mathcal{H}_{n})=\text{dim}(\mathcal{H})$ for all $n$.
+If $\text{dim}(H_{n})=\text{dim}(H)$ for all $n$, then $\text{dim}(\mathcal{H}_{n})=\text{dim}(\mathcal{H})$ for all $n$.
 By Lemma 2.1, $\text{dim}(V)=\text{dim}(\mathcal{H})$.
@@
 Thus, $\mathcal{H}_{n}\to\mathcal{H}$ in SubG.
-Assume that $\exp\colon\mathcal{G}\to G$ is a diffeomorphism and $H_{n}$ are connected.
-Then $\exp\colon\mathcal{H}_{n}\to H$ is also a diffeomorphism for all $n$.
-Let $h\in H$.
-Since $H_{n}\to H$, there are $h_{n}\in H_{n}$ such that $h_{n}\to h$.
-Let $X_{n}\in\mathcal{H}_{n}$ and $X\in\mathcal{H}$ be such that $\exp(X_{n})=h_{n}$ and $\exp(X)=h$.
-Since $\exp$ is a diffeomorphism, we get that $X_{n}\to X$.
-Since $\mathcal{H}_{n}\to V\subset\mathcal{H}$, we get that $X\in\mathcal{H}$.
-Thus, $\exp(\mathcal{H})=H$.
-Therefore, $H$ is connected.
-Since $\exp$ is a diffeomorphism, we also get that $\mathcal{H}_{n}\to\mathcal{H}$ and hence $V=\mathcal{H}$.
-Thus, ${\rm limsup}~{}\text{dim}(H_{n})=\text{dim}(H)$.
+3. On the support of random subgroups invariant under unipotent transformations
+In this section we consider invariant measures on Grassmannian for projective linear actions and first prove a result on the support of the invariant measure.
 The result of the following kind is known as Furstenberg’s Lemma (cf.
@@
 Then $I_{\lambda}$ is a normal algebraic subgroup of $G_{\lambda}$.
-Since ${\rm Gr}_{k}(V)\simeq GL(V)/H$ where $H$ is the stablizer of a $k$-dimensional subspace, by Corollary 2.6 of [5] we get that $G_{\lambda}$ is a real algebraic group and $G_{\lambda}/I_{\lambda}$ is compact.
+Since ${\rm Gr}_{k}(V)\simeq GL(V)/H$ where $H$ is the stabilizer of a $k$-dimensional subspace, by Corollary 2.6 of [7] we get that $G_{\lambda}$ is a real algebraic group and $G_{\lambda}/I_{\lambda}$ is compact.
 Thus, any unipotent transformation in $G_{\lambda}$ is in $I_{\lambda}$.
@@
 ∎
+The following proposition is crucial to the proof of our results, it describes the action of the unipotent group on orbits that support invariant measures: for a connected Lie group $G$, Aut(G)(G) has a natural action on the Lie algebra of $G$ by identifying each $\tau\in{\rm Aut}(G)$ with its differential ${\rm d}\tau$.
 Proposition 3.2.
-Let $G$ be a connected Lie group with a closed connected Lie subgroup $S$.
-Let $\Gamma$ be a $\sigma$-compact subgroup of Aut$(G)$ and $U$ be a subgroup of $\Gamma$ consisting of unipotent automorphisms.
-Suppose there is a ${\Gamma}$-invariant probability measure $\mu$ on ${\rm Sub}_{G}$ such that $\mu(\{\tau(S)\mid\tau\in\Gamma\})=1$.
-Then $S$ is $N$-invariant where $N$ is the smallest normal subgroup of $\Gamma$ containing $U$.
+Let $G$ be a connected Lie group with Lie algebra $\mathcal{G}$}, and $S$ be a closed connected subgroup of $G$ with Lie algebra $\mathcal{S}$.
+Let $\Gamma$ be a $\sigma$-compact subgroup of Aut(G)(G), and $U$ be a subgroup of $\Gamma$ consisting of unipotent automorphisms.
+If either
+(1) there is a $\Gamma$}-invariant probability measure $\lambda$ on ${\rm Sub}_{\mathcal{G}}$ with $\lambda(\{{\rm d}\tau({\mathcal{S}})\mid\tau\in\Gamma\})=1$
+or
+(2) there is a $\Gamma$}-invariant probability measure $\mu$ on ${\rm Sub}_{G}$ with $\mu(\{\tau(S)\mid\tau\in\Gamma\})=1$;
+then $S$ is $N$-invariant where $N$ is the smallest normal subgroup of $\Gamma$ containing $U$
 Proof.
-Let $\mathcal{G}$ and $\mathcal{S}$ be the Lie algebras of $G$ and $S$ respectively.
-Let $\mathcal{O}=\{\tau(S)\mid\tau\in{\rm Aut}~{}(G)\}$.
 Let $X={\rm Gr}_{k}(\mathcal{G})$ be the Grassmannian manifold of $k$-dimensional subspaces of $\mathcal{G}$ where $k$ is the dimension of $\mathcal{S}$.
-Let $\phi\colon\mathcal{O}\to X$ be defined by $\phi(\tau(S))={\rm d}\tau(\mathcal{S})$.
+Consider the canonical action of $\Gamma$ on $X$ given by: $(\tau,V)\mapsto{\rm d}\tau(V)$ for $\tau\in\Gamma$ and $V\in X$.
+If (1) holds, then in view of Lemma 2.1, we get that $\lambda$ is a $\Gamma$-invariant probability measure on $X$.
+By Lemma 3.1, $\lambda$ is supported on $U$-invariant subspaces.
+This proves that $\mathcal{S}$ is NN-invariant.
+Since $S$ is a connected Lie subgroup, $S$ is NN-nvariant.
+Suppose (2) holds.
+Let $\mathcal{O}=\{\tau(S)\mid\tau\in{\rm Aut}~(G)\}$ and $\phi\colon\mathcal{O}\to X$ be defined by $\phi(\tau(S))={\rm d}\tau(\mathcal{S})$.
 Then it follows from Proposition 2.2 and Lemma 2.1 that $\phi$ is continuous.
 Let $\lambda=\phi(\mu)$.
-Consider the canonical action of $\Gamma$ on $X$ given by: $(\tau,V)\mapsto{\rm d}\tau(V)$ for $\tau\in\Gamma$ and $V\in X$.
-Then $\lambda$ is a $\Gamma$-invariant probability measure on $X$.
-By Lemma 2.4, $\lambda$ is supported on $U$-invariant subspaces.
-This proves that $\mathcal{S}$ is $N$-invariant.
-Since $S$ is a connected Lie subgroup, $S$ is $N$-nvariant.
+Then $\lambda$ is a $\Gamma$-invariant probability measure and $\lambda(\{\tau({\mathcal{S}})\mid\tau\in\Gamma\})=1$.
+Now the result follows from the previous case.
 ∎
 4. IRS on orbits
-In this section, we show the existence and non-existence of invariant random subgroups supported on various orbits of the conjugation action of a connected Lie group $G$ on SubG.
+In this section, we show the existence of invariant random subgroups supported on various orbits of the conjugation action of a connected Lie group $G$ on $\text{Sub}_{G}$.
 In other words, we discuss the proofs of Theorems 1.1, 1.3, 1.4 and 1.6.
@@
 Let $\mathcal{G}$ be the Lie algebra of $G$.
-Since $N$ is a nilpotent normal subgroup of $G$, Ad$(x)$ is a unipotent transformation on $\mathcal{G}$ for all $x\in N$.
+Since $N$ is a nilpotent normal subgroup of $G$, Ad(x)(x) is a unipotent transformation on $\mathcal{G}$ for all $x\in N$.
 Therefore, by Proposition 3.2, $\mu$ is supported on $N$-invariant subspaces of $\mathcal{G}$.
@@
 Since any two semisimple Levi subgroups are conjugate by an element of $N$, we get that $L$ is normal in $G$.
-Suppose the set of all semisimple Levi subgroups is closed in SubG.
+Suppose the set of all semisimple Levi subgroups is closed in $\text{Sub}_{G}$.
 Then $L_{G}$ is compact; and, since $N$ is amenable, $L_{G}$ admits an $N$-invariant probability measure.
@@
 Now, let $N$ be the nilpotent radical of $G$.
-Then Ad$(x)$ is unipotent for any $x\in N$, hence $N$ normalizes $K$.
+Then Ad(x)(x) is unipotent for any $x\in N$, hence $N$ normalizes $K$.
 Thus, $NK$ is a closed connected subgroup of $G$.
@@
 Proposition 4.1.
-Let $G$ be a connected Lie group with solvable radical $R$ and $M$ be a maximal compact subgroup of $G$. Suppose $M$ is normal in $G$. Then we have the following:
+Let $G$ be a connected Lie group with solvable radical $R$ and $M$ be a maximal compact subgroup of $G$. Suppose $M$ is normal in $G$. Then,
   1. (1)
-    solvable radical $R$ admits a compact central subgroup $T$ so that $R/T$ is simply connected;
+    the solvable radical $R$ admits a compact central subgroup $T$ so that $R/T$ is simply connected;
   2. (2)
@@
 ∎
+We now apply the above results to show that a finite covolume subgroup is cocompact if the subgroup is a Levi subgroup or a maximal compact subgroup.
+Corollary 4.3.
+Let $G$ be a connected Lie group and $H$ be a closed subgroup of $G$ such that $G/H$ carries a $G$-invariant probability measure.
+Suppose $H$ is either a Levi subgroup or a maximal compact subgroup.
+Then $G/H$ is a compact group.
+Proof.
+We prove the result when $H$ is a Levi subgroup, the other case is similar.
+Suppose $H$ is a Levi subgroup of $G$.
+Then the stabilizer, say $G_{xH}$, at any $xH\in G/H$ is $xHx^{-1}$.
+Therefore, using the map $xH\mapsto G_{xH}$, we obtain a IRS supported on the set of semisimple Levi subgroups.
+By Theorem 1.1, we get that $G$ has a unique semisimple Levi subgroup and is normal in $G$.
+This implies that $G/H$ is a locally compact group.
+Since $G/H$ carries a $G$-invariant probability measure, $G/H$ is compact.
+∎
 Now, we assume that $G$ is a real algebraic group. Also, let $B$ be a Borel subgroup of $G$ and $B_{G}$ be the set of all Borel subgroups of $G$, i.e., $B_{G}=\{xBx^{-1}|x\in G\}$.
@@
 Since any two Borel subgroups are conjugate in $G$, $B_{G}=\{xBx^{-1}\mid x\in C\}$.
-Thus, $B_{G}$ is a compact subset of SubG.
+Thus, $B_{G}$ is a compact subset of $\text{Sub}_{G}$.
 Suppose that there is an IRS $\mu$ such that $\mu(B_{G})=1$.
@@
 Suppose that there is an IRS $\mu$ such that $\mu(D_{G})=1$.
-Then, by Proposition 2.5, $D$ is normalized by any unipotent subgroup of $G$.
-If $S$ is a noncompact simple algebraic subgroup of $G$, then maximal split diagonalizable subgroup $D_{S}$ of $S$ is contained in a conjugate of $D$.
+Then, by Proposition 3.2, Any unipotent subgroup of $G$ normalizes $D$.
+If $S$ is a noncompact simple algebraic subgroup of $G$, then the maximal split diagonalizable subgroup $D_{S}$ of $S$ is contained in a conjugate of $D$.
 Replacing $S$ by a conjugate of $S$, we may assume that $D_{S}\subset D$.
@@
 This implies that $D\subset Z(L)$ because $L=[L,L]Z(L)$.
-Since $D$ is normalized by the unipotent subgroups, $D$ is normalized by $U$. Since $D\cap U$ is trivial, $D$ is centralized by $U$ also. Since $G=LU$, $D$ is central in $G$.
-The maximal diagonalizable subgroup being normal has certain implications on the structure of $G$ which is noted down below.
-Proposition 3.3.
-Let $G$ be a connected real algebraic group and $D$ be a maximal diagonalizable subgroup of $G$.
-In addition, suppose $D$ is normal in $G$.
-Then any Levi subgroup of $G$ is a compact subgroup, and the solvable radical is a compact extension of a nilpotent group.
-Proof.
-Let $H$ be a noncompact simple factor of a semisimple Levi subgroup of $G$ and $D_{H}$ be the maximal diagonalizable subgroup of $H$.
-Then $D_{H}$ is contained in a conjugate of $D$.
-Since $D$ is normal, $D_{H}\subset D$.
-This implies that $D\cap H$ is an abelian normal subgroup of positive dimension in $H$, which is a contradiction to the simplicity of $H$.
-Therefore, any semisimple Levi subgroup of $G$ is compact.
-Now, let $R$ be the solvable radical of $G$.
-Then $R\cap D$ is a normal subgroup $R$, hence $D_{R}U$ is a nilpotent subgroup of $R$ where $D_{R}=D\cap R$ and $U$ is the unipotent radical of $G$.
-Since $R/D_{R}U$ is a anisotropic torus (that is, without any split torus subgroup) we get that $R/D_{R}U$ is compact.
-∎
-4. Application to measure preserving actions
-Let $G$ be a locally compact group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations.
-It follows from [14] that the stabilizer of almost every point in X is closed in G and the stabilizer map $x\mapsto G_{x}$ is measurable where $G_{x}=\{g\in G\mid gx=x\}$ for $x\in X$.
-Thus, the $G$-invariant probability measure $m$ on $X$ may be realized as an IRS on G which leads to the following:
-Corollary 4.1.
-Let $G$ be a connected Lie group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations.
-Then almost none of the stabilizer subgroups is a semisimple Levi subgroup (resp.
-maximal compact subgroup) of $G$ or $m$ has an atom $x$ such that $G_{x}$ is a semisimple Levi subgroup (resp.
-maximal compact subgroup) of $G$ and $G_{x}$ is normal in $G$.
-In addition, if the action of $G$ is ergodic, then almost none of the stabilizer is a semisimple Levi subgroup of $G$ or the action of $G$ factors through the solvable radical of $G$.
-Proof.
-Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$ and $X_{L}=\{x\in X\mid G_{x}\in L_{G}\}$.
-Since $L_{G}$ is invariant under the conjugation action and $G_{gx}=gG_{x}g^{-1}$, it follows that $X_{L}$ is invariant under the action of $G$.
-Suppose $m(X_{L})>0$.
-Then, replacing $X$ by $X_{L}$ and $m$ by the restriction of $m$ to $X_{L}$, we obtain a measure preserving action of $G$ such that almost all stabilizers are Levi subgroups of $G$.
-Using the identification of $x\mapsto G_{x}$, we obtain a IRS supported on the set of semisimple Levi subgroups.
-By Theorem 1.1, we get that $G$ has a unique semisimple Levi subgroup, say L, and $L$ is normal in $G$.
-The case of maximal compact subgroups may be proved in a similar way.
-Suppose the action of $G$ is ergodic.
-Then $X_{L}$ has full measure or null.
-In case $X_{L}$ has full measure, any $G_{x}=G_{y}$ for all $x,y\in X_{L}$ and hence the semisimple Levi subgroup acts trivially on $X$, hence the action factors through the solvable radical of $G$.
-A similar application can also be done for maximal diagonalizable groups and Borel subgroups as in the following: we omit the details as it is same as in Corollary 4.1
-Corollary 4.2.
-Let $G$ be a connected real algebraic group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations.
-Then almost none of the stabilizer subgroup is a Borel (resp.
-maximal diagonalizable) subgroup of $G$ or $m$ has an atom $x$ such that $G_{x}$ is a Borel (resp.
-maximal diagonalizable) subgroup of $G$ and $G_{x}$ is normal in $G$.
-In addition if the action of $G$ is ergodic, then almost none of the stabilizer is a Borel (resp.
-maximal diagonalizable) subgroup of $G$ or the action of $G$ factors through a compact subgroup (resp.
-compact extension of the unipotent radical) of $G$.
-Corollary 4.3.
-Let $G$ be a connected Lie group and $K$ be a maximal compact subgroup of $G$. Suppose $G/K$ carries a $G$-invariant probability measure. Then $G$ is a compact group.
-Proof.
-Let $m$ be the $G$-invariant probability measure on $G/K$.
-Then $(G/K,m)$ is a probability space on which $G$ acts by measure preserving transformations and stabilizer at any $xK\in G/K$ is $xKx^{-1}$.
-Therefore by Corollary 4.1, $K$ is normal in $G$, hence $G/K$ is a locally compact group.
-Since $G/K$ carries a $G$-invariant probability measure, $G/K$ is compact.
-Thus, $G$ is compact.
+Since the unipotent subgroups normalize $D$, $D$ is normalized by $U$.
+Since $D\cap U$ is trivial, $D$ is centralized by $U$ also.
+Since $G=LU$, $D$ is central in $G$.
+This proves that (1) implies (2) and that (2) implies (3) is trivial.
+If $D$ is normal, the $\delta_{D}$, the dirac measure at $D$ is the required IRS.
+This proves that (3) implies (1).
+Suppose (1) holds.
+Then the proof of the first implication shows that Levi subgroup is compact and $D$ is central.
+Let $R$ be the solvable radical of $G$.
+Then $D_{R}=R\cap D$ is a central subgroup of $R$, hence $D_{R}U$ is a nilpotent subgroup of $R$ where $U$ is the unipotent radical of $G$.
+Since $R/D_{R}U$ is an anisotropic torus (that is, without any split torus subgroup) we get that $R/D_{R}U$ is compact.
 ∎
@@
 Let $m$ be the $G$-invariant probability measure on $G/D$.
-Then $(G/D,m)$ is a probability space on which $G$ acts by measure preserving transformations and stabilizer at any $xD\in G/D$ is $xDx^{-1}$.
-Therefore by Corollary 4.2, $D$ is normal in $G$.
-By Proposition 3.3, the Levi subgroup of $G$ is a compact normal subgroup and the solvable radical is nilpotent.
-Since $G/D$ has a $G$-invariant probability measure, unipotent radical of $G$ is trivial.
-Thus, $G$ is a direct product of $D$ and a compact group.
+Then stabilizer, $G_{xD}$, at any $xD\in G/D$ is $xDx^{-1}$.
+Therefore, the map $xD\mapsto xDx^{-1}$ gives rise to an IRS supported on maximal diagonalizable subgroups.
+By Theorem 1.6, we get that $D$ is normal in $G$ and the Levi subgroup of $G$ is compact.
+Since $G/D$ has a $G$-invariant probability measure, $G/D$ is a compact group.
+Therefore, the unipotent radical of $G$ is trivial.
+Thus, $G$ is a direct product of $D$ and a compact subgroup.
 ∎
-5. Conjugacy limits in $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$
-5.1. Limits of levi subgroups
-From now onwards, we use the notation $G$ for $\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$, and the notation $L$ for the group $\operatorname{SL}(2,\mathbb{R})$ inside $G$.
-Also, let $L_{G}$ be the set of all semsimple Levi subgroups inside $G$.
-Then $L_{G}$ consists of conjugates of $L$ by elements of $G$.
-We denote elements in $G$ by $(g,v)$ with $g\in\operatorname{SL}(2,\mathbb{R})$ and $v\in\mathbb{R}^{2}$.
-Now, if $(G_{n})$ is a sequence of Levi subgroups in $L_{G}$, then each $G_{n}$ is given by a conjugate of $L$ by some elements of $G$, i.e.,
-$G_{n}=\ (g_{n},v_{n})\ L\ (g_{n},v_{n})^{-1},$
-where $(g_{n},v_{n})^{-1}$ denotes the inverse of $(g_{n},v_{n})$ in $G$.
-If we conjugate $L$ by elements of the form $(g,(0,0))$, the resultant is $L$ itself.
-Also, note that $(g,v)=(I,v)\ (g,(0,0))$, where $I$ is the identity element in $\operatorname{SL}(2,\mathbb{R})$.
-Therefore, to determine all the limit points of $L_{G}$, it is enough to consider the conjugacy limits of the form
-$(I,v_{n})\ L\ (I,v_{n})^{-1}.$
-The following lemma is a key ingredient for the results in this section and afterwards.
-Lemma 5.1.
-Let $(g_{n})$ and $(v_{n})$ be sequences in $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ respectively such that $v_{n}\not\to 0$ and
-$(I,v_{n})\ (g_{n},(0,0))\ (I,v_{n})^{-1}=(g_{n},v_{n}-g_{n}(v_{n}))\rightarrow(g,(0,0))$
-for some $g\in\operatorname{SL}(2,\mathbb{R})$. Then $g$ is unipotent.
-Proof.
-Suppose $(g_{n})$ and $(v_{n})$ are sequences in $\operatorname{SL}(2,\mathbb{R})$ and $\mathbb{R}^{2}$ respectively such that $v_{n}\not\to 0$ and
-$(I,v_{n})\ (g_{n},(0,0))\ (I,v_{n})^{-1}=(g_{n},v_{n}-g_{n}(v_{n}))\rightarrow(g,(0,0))$
-for some $g\in\operatorname{SL}(2,\mathbb{R})$.
-Then $g_{n}\to g$ in $SL_{2}(\mathbb{R})$ and $(I-g_{n})(v_{n})\rightarrow(0,0)$ as $n\rightarrow\infty$.
-Note that if $I-g$ is invertible, then since $g_{n}\to g$, $I-g_{n}\to I-g$.
-Since invertible matrices form an open set, $I-g_{n}$ is invertible for all large $n$ and hence $(I-g_{n})^{-1}\to(I-g)^{-1}$.
-As $(I-g_{n})(v_{n})\rightarrow(0,0)$, we arrive at a contradiction to the assumption that $v_{n}\not\to 0$.
-Therefore, $I-g$ is not invertible.
-Since $g\in SL(2,\mathbb{R})$, $g$ is an unipotent transformation.
-∎
-Proof of Theorem 1.7.
-Assume that $(I,v_{n})\ L\ (I,v_{n})^{-1}$ converges to $H$ in $\text{Sub}_{G}$; we want to determine $H$. The proof is divided into following caes.
-Case $1$.
-Suppose $v_{n}\rightarrow v$ as $n\rightarrow\infty$.
-In this case, it is easy to see that $(I,v_{n})\ L\ (I,v_{n})^{-1}$ converges to the closed subgroup $(I,v)\ L\ (I,v)^{-1}$, and hence the limit $H$ is a Levi subgroup.
-We now assume that $(v_{n})$ is unbounded. It follows from Lemma 5.1 that if $g\in SL(2,\mathbb{R})\cap H$, then $g$ is unipotent.
-Case $2$.
-Suppose $v_{n}=(\alpha_{n},\beta_{n})$ and $\alpha_{n}$ and $\beta_{n}$ both diverge.
-We first show that the subgroup $\mathbb{R}^{2}$ is contained in $H$.
-Let $\mathfrak{g}_{n}=\begin{pmatrix}1&\frac{-s}{\beta_{n}}\\ 0&1\end{pmatrix}$.
-Then
-$(I,v_{n})\ (\mathfrak{g}_{n},(0,0))\ (I,v_{n})^{-1}=(\mathfrak{g}_{n},v_{n}-\mathfrak{g}_{n}(v_{n}))=(\mathfrak{g}_{n},(s,0))\rightarrow(I,(s,0))$
-as $n\rightarrow\infty$. Similarly, considering
-$\mathfrak{h}_{n}=\begin{pmatrix}1&0\\ \frac{-t}{\alpha_{n}}&1\end{pmatrix}$
-it is easy to see that
-$(I,v_{n})\ (\mathfrak{h}_{n},(0,0))\ (I,v_{n})^{-1}\rightarrow(0,t)$
-as $n\rightarrow\infty$. Then composing $(I,(s,0))$ and $(I,(0,t))$, we see that $(I,(s,t))\in H$ for any $s,t\in\mathbb{R}$.
-Since $G=SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ and $\mathbb{R}^{2}\subset H$, we get that $H=(H\cap L)\ltimes\mathbb{R}^{2}$.
-But $L\cap H$ consists of unipotent transformations.
-Since $N^{+}$ and its conjugates in $SL(2,\mathbb{R})$ are the maximal subgroups with unipotent elements, $H$ is given by
-$H=hN^{+}h^{-1}\ltimes\mathbb{R}^{2}$
-for some $h\in\operatorname{SL}(2,\mathbb{R})$, where
-$N^{+}=\left\{\begin{pmatrix}1&s\\ 0&1\end{pmatrix}:s\in\mathbb{R}\right\},$
-Case $3$.
-The sequence $(v_{n})=(\alpha_{n},\beta_{n})$ diverges but one of $\alpha_{n}$ or $\beta_{n}$ is bounded.
-First suppose, $v_{n}=(\alpha_{n},0)$ with $\alpha_{n}$ diverging.
-Then, for any real number $s$, if we take $\mathfrak{g}_{n}:=\begin{pmatrix}\frac{\displaystyle{\alpha_{n}-s}}{\displaystyle{\alpha_{n}}}&0\\ 0&\frac{\displaystyle{\alpha_{n}}}{\displaystyle{\alpha_{n}-s}}\end{pmatrix}$, we see that
-$(I,v_{n})(\mathfrak{g}_{n},(0,0))(I,v_{n})^{-1}\to(I,(s,0)).$
-Similarly, considering $\mathfrak{h}_{n}:=\begin{pmatrix}1&0\\ \frac{\displaystyle{-s}}{\displaystyle{\alpha_{n}}}&1\end{pmatrix}$, we see that
-$(I,v_{n})(\mathfrak{h}_{n},(0,0))(I,v_{n})^{-1}\to(0,s).$
-This shows that $\mathbb{R}^{2}$ is contained in the limit subgroup $H$.
-Also, since any element of $N^{+}$ fixes $v_{n}$, it follows that $N^{+}$ is contained in $H$ as well.
-Now it follows, as in Case $2$, that
-$H=N^{+}\ltimes\mathbb{R}^{2}.$
-Similarly, if $v_{n}=(0,\alpha_{n})$ with $\alpha_{n}$ diverging, then
-$H=N^{-}\ltimes\mathbb{R}^{2}.$
-where
-$N^{-}=\left\{\begin{pmatrix}1&0\\ s&1\end{pmatrix}:s\in\mathbb{R}\right\}.$
-Now suppose $v_{n}=(\alpha_{n},\beta_{n})$ with $\alpha_{n}$ diverging and $\beta_{n}$ bounded.
-As we are looking for limit points of the set of all Levi subgroups, we may assume in this case that $\alpha_{n}$ diverges and $\beta_{n}$ converges to some $\beta$.
-We can write $(I,(\alpha_{n},\beta_{n}))=(I,(\alpha_{n},0))(I,(0,\beta_{n}))$.
-So, if $(I,(\alpha_{n},0))L(I,(\alpha_{n},0))^{-1}$ converges to the subgroup $H$, then
-$(I,v_{n})L(I,v_{n})^{-1}$ converges to the subgroup $(I,(0,\beta))H(I,(0,\beta))^{-1}$.
-Therefore, $H$ is a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
-Thus, any subgroup in $\overline{L_{G}}$ is either a Levi subgroup or a conjugate of $N^{+}\ltimes\mathbb{R}^{2}$.
-5.2. Limits of maximal compact subgroups
-We now identify the limit points of the set of maximal compact subgroups. In other words, we find the closure of the orbit of $K$ under the conjugacy action of $G$ inside $\text{Sub}_{G}$.
-Proof of Theorem 1.8.
-Let us recall that $K$ is given by
-$K:=\ \left\{\begin{pmatrix}\text{cos}\ \theta&\text{sin}\ \theta\\ -\text{sin}\ \theta&\text{cos}\ \theta\end{pmatrix}\ :\ \theta\in\mathbb{R}\right\}.$
-In order to find the limit points of
-$K_{G}=\{(g,v)K(g,v)^{-1}\mid g\in SL(2,\mathbb{R}),v\in\mathbb{R}^{2}\},$
-it is sufficient to identify the limit points up to conjugation.
-Using Iwasawa decomposition, i.e., $SL(2,\mathbb{R})=N^{+}AK$, we may assume that $g\in N^{+}A$.
-Let $H$ be a limit point of $K_{G}$.
-Then there exist sequences
-$\mathfrak{a}_{n}=\begin{pmatrix}a_{n}&0\\ 0&a_{n}^{-1}\end{pmatrix}\in A,~{}\mathfrak{s}_{n}=\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix}\in N^{+}$
-and $v_{n}=(\alpha_{n},\beta_{n})\in\mathbb{R}^{2}$ such that $(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})K(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}\to H$.
-By passing to a subsequence and replacing $\mathfrak{a}_{n}$ or $\mathfrak{s}_{n}$ by their conjugates, we may assume that $a_{n}\to\infty$ or $a_{n}=1$ and $s_{n}\to\infty$ or $s_{n}=0$.
-Let $(g,v)\in H$. Then there exists
-$\Theta_{n}=\begin{pmatrix}\cos\theta_{n}&\sin\theta_{n}\\ -\sin\theta_{n}&\cos\theta_{n}\end{pmatrix}\in K$
-such that $(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})\Theta_{n}(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}\to(g,v)$
-Now,
-$(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})\Theta_{n}(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}=(T_{n},(x_{n},y_{n})),$
-where
-$T_{n}=\begin{pmatrix}\cos\theta_{n}-\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}&(\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\\ -\sin\theta_{n}/a_{n}^{2}&\cos\theta_{n}+\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\end{pmatrix}$
-and
-$x_{n}=\alpha_{n}\big{(}1-\cos\theta_{n}+{\frac{s_{n}}{a_{n}^{2}}}\sin\theta_{n}\big{)}+\beta_{n}\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n},$
-$y_{n}=-\alpha_{n}\frac{\sin\theta_{n}}{a_{n}^{2}}+\beta_{n}\big{(}1-\cos\theta_{n}-\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\big{)}.$
-We now assume that $v_{n}\to\infty$. As $\big{(}T_{n},(x_{n},y_{n})\big{)}\to(g,v)$, it follows from Lemma 5.1 that $g$ is unipotent.
-Case $1$: We now consider the case when ${\frac{\alpha_{n}}{\beta_{n}}}\to 0$. Since $v_{n}$ diverges, $\beta_{n}\to\infty$. Let $x,y\in\mathbb{R}$ be such that $v=(x,y)$. Then
-$x_{n}=\alpha_{n}\big{(}1-\cos\theta_{n}+\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\big{)}+\beta_{n}\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n}\to x.$
-Since $T_{n}\to g$, $\cos\theta_{n}-\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}$ converges. Therefore,
-$x_{n}=\beta_{n}\Big{[}\frac{\alpha_{n}}{\beta_{n}}\big{(}1-\cos\theta_{n}+\frac{s_{n}}{a_{n}^{2}}\sin\theta_{n}\big{)}+\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n}\Big{]}$
-converges implies that $\big{(}\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2}\big{)}\sin\theta_{n}\to 0$. Since $g$ is unipotent, it follows that $g=I$.
-Also, for any $t\in\mathbb{R}$, we can choose $\zeta_{n}\to 0$ so that $\beta_{n}(\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2})\sin\zeta_{n}\to t$.
-Then $\beta_{n}\frac{s_{n}}{a_{n}^{2}}\sin\zeta_{n}\to 0$ and $\beta_{n}\sin\zeta_{n}$ converges (to 00 or $t$ depending on if $s_{n}^{2}+a_{n}^{2}\to\infty$ or not).
-Since $\beta_{n}\to\infty$, we also get that $(\frac{s_{n}^{2}}{a_{n}^{2}}+a_{n}^{2})\sin\zeta_{n}\to 0$.
-Since $\zeta_{n}\to 0$ and $\beta_{n}\to\infty$, we get that $\mathfrak{s}_{n}\mathfrak{a}_{n}Z_{n}\mathfrak{a}_{n}^{-1}\mathfrak{s}_{n}^{-1}\to I$ for $Z_{n}=\begin{pmatrix}\cos\zeta_{n}&\sin\zeta_{n}\\ -\sin\zeta_{n}&\cos\zeta_{n}\end{pmatrix}\in K$.
-Since $\zeta_{n}\to 0$, $(1-\cos\zeta_{n})/\sin\zeta_{n}\to 0$, hence $\alpha_{n}(1-\cos\zeta_{n})=(\beta_{n}\sin\zeta_{n})(\frac{\alpha_{n}}{\beta_{n}})(\frac{1-\cos\zeta_{n}}{\sin\zeta_{n}})\to 0$ and $-\alpha_{n}{\sin\zeta_{n}\over a_{n}^{2}}+\beta_{n}\big{(}1-\cos\zeta_{n}-{s_{n}\over a_{n}^{2}}\sin\zeta_{n}\big{)}={-\alpha_{n}\over\beta_{n}}{\beta_{n}\sin\zeta_{n}\over a_{n}^{2}}+\beta_{n}\sin\zeta_{n}({1-\cos\zeta_{n}\over\sin\zeta_{n}})-\beta_{n}{s_{n}\over a_{n}^{2}}\sin\zeta_{n}\to 0$.
-Thus,
-$(\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})\ Z_{n}\ (\mathfrak{s}_{n}\mathfrak{a}_{n},v_{n})^{-1}\to(I,(t,0)).$
-Now, replacing $\Theta_{n}$ by $\Theta_{n}Z_{n}^{-1}$, we may assume that $x_{n}\to 0$.
-Since $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\to 0$ and ${s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2}\not\to 0$, we get that $\sin\theta_{n}\to 0$.
-Replacing $\theta_{n}$ by a translate of $2k\pi$, we may assume that $\theta_{n}\to 0$.
-This implies that ${1-\cos\theta_{n}\ \over\sin\theta_{n}}\to 0$.
-Since $x_{n}\to 0$, we get that $\beta_{n}({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\to 0$, hence $\alpha_{n}(1-\cos\theta_{n}+{s_{n}\over a_{n}^{2}}\sin\theta_{n})\to 0$.
-Therefore,
-$|\alpha_{n}{-\sin\theta_{n}\over a_{n}^{2}}|\leq|{\alpha_{n}\over\beta_{n}}||\beta_{n}({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}|\to 0$
-and
-$\begin{array}[]{cl}&|\beta_{n}(1-\cos\theta_{n}-{s_{n}\over a_{n}^{2}}\sin\theta_{n})|\\ =&|{1-\cos\theta_{n}\over\sin\theta_{n}}{a_{n}^{2}\over s_{n}^{2}+a_{n}^{4}}-{s_{n}\over s_{n}^{2}+a_{n}^{4}}||\beta_{n}({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}|\to 0.\end{array}$
-This implies that $H=V_{0}=\{(I,(x,0))\mid x\in\mathbb{R}\}$.
-Case $2$: Suppose ${\beta_{n}\over\alpha_{n}}\to 0$.
-Let $F=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$.
-Since $(g_{n},v_{n})K(g_{n},v_{n})^{-1}\to H$, conjugating by $F$, we get that $(Fg_{n},Fv_{n})K(Fg_{n},Fv_{n})^{-1}\to FHF^{-1}$.
-Since $Fv$ satisfies case (i) we get that $FHF^{-1}=V_{0}$.
-Thus, $H=V_{\infty}=\{(I,(0,y)\mid y\in\mathbb{R}\}$.
-Case $3$: Suppose $c\not=0$ is a limit point of $|{\alpha_{n}\over\beta_{n}}|$.
-By passing to a subsequence we may assume that ${\alpha_{n}\over\beta_{n}}\to c$.
-There are two cases here: if one of $a_{n}$ or $s_{n}$ diverges to $\infty$, then as in case (i) we can show that $H=V_{0}$.
-We now assume that $a_{n}=1$ and $s_{n}=0$.
-Then, choosing $\zeta_{n}\rightarrow 0$ such that $\beta_{n}\sin\zeta_{n}\rightarrow t$ for any preassigned real number $t$, it is easy to see that
-$g_{n}Z_{n}g_{n}^{-1}\rightarrow\big{(}I,\ (t,-ct)\big{)}.$
-for suitable $Z_{n}\in K$. Therefore, in this case as in case (i) we may show that $H=V_{c}=\{(I,(t,ct))\mid t\in\mathbb{R}\}$.
-We now consider the case when $(v_{n})$ has a convergent subsequence.
-In this case, by passing to a subsequence and conjugating by $(I,-v_{n})$, we may assume that $(\mathfrak{s}_{n}\mathfrak{a}_{n},0)K(\mathfrak{s}_{n}\mathfrak{a}_{n},0)^{-1}\to H$.
-Since $\mathfrak{s}_{n},\mathfrak{a}_{n}\in SL(2,\mathbb{R})$ and $K\subset SL(2,\mathbb{R})$, $H\subset SL(2,\mathbb{R})$.
-Recall that we also have $s_{n}\to\infty$ or $s_{n}=0$ and $a_{n}\to\infty$ or $a_{n}=1$.
-If both $s_{n}=0$ and $a_{n}=1$, we are done.
-So we may assume that either $s_{n}\to\infty$ or $a_{n}\to\infty$.
-We consider the case that $s_{n}\to\infty$ and other case that $a_{n}\to\infty$ is similar.
-We have
-$T_{n}=\begin{pmatrix}\cos\theta_{n}-{s_{n}\over a_{n}^{2}}\sin\theta_{n}&({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\\ {-\sin\theta_{n}/a_{n}^{2}}&\cos\theta_{n}+{s_{n}\over a_{n}^{2}}\sin\theta_{n}\end{pmatrix}\to g=\begin{pmatrix}g_{1,1}&g_{1,2}\\ g_{2,1}&g_{2,2}\end{pmatrix},{~{}~{}\rm say}.$
-This in particular, implies that $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}$ converges.
-Since ${s_{n}\over a_{n}^{2}}+a_{n}^{2}\to\infty$, $\sin\theta_{n}\to 0$ and since ${\sin\theta_{n}\over a_{n}^{2}}\to g_{2,1}$, we get that $g_{2,1}=0$.
-Also, for $t\in\mathbb{R}$, we choose $\zeta_{n}\to 0$ such that $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\zeta_{n}\to t$.
-Then $\cos\zeta_{n}\to 1$ and ${s_{n}\over a_{n}^{2}}\sin\zeta_{n}\to 0$, hence $g_{n}Z_{n}g_{n}^{-1}\to\begin{pmatrix}1&t\\ 0&1\end{pmatrix}$ for suitable $Z_{n}\in K$.
-If, in addition $g_{1,2}=0$, then $({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\sin\theta_{n}\to 0$ and hence ${s_{n}\over a_{n}^{2}}\sin\theta_{n}\to 0$.
-Thus, $\cos\theta_{n}$ converges to $g_{1,1}$ as well as $g_{2,2}$ but $g_{1,1}g_{2,2}=1$.
-Therefore, $g_{1,1}=\pm 1=g_{2,2}$.
-In fact, taking $\theta_{n}=\pi$, we get that
-$g_{n}\Theta_{n}g_{n}^{-1}=\begin{pmatrix}\cos\theta_{n}-{s_{n}\over a_{n}^{2}}\ \sin\theta_{n}&({s_{n}^{2}\over a_{n}^{2}}+a_{n}^{2})\ \sin\theta_{n}\\ -\sin\theta_{n}/a_{n}^{2}&\cos\theta_{n}+{s_{n}\over a_{n}^{2}}\sin\theta_{n}\end{pmatrix}=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}.$
-Thus, we have $H=\tilde{N^{+}}$, that is $g_{n}Kg_{n}^{-1}\rightarrow\tilde{N^{+}}$.
-5.3. Limits of maximal diagonalizable subgroups
-Now we study limit points of the set of maximal diagonalizable subgroups of $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-We consider the conjugacy limits $(g_{n},v_{n})A(g_{n},v_{n})^{-1}$ of $A$ for different types of $g_{n}$ and $v_{n}$.
-Proof of Theorem 1.9.
-Let $A=\left\{\begin{pmatrix}a&0\\ 0&1/a\end{pmatrix}\mid a\in\mathbb{R}^{+}\right\}$ and $H$ be a limit point of $\{(g,v)A(g,v)^{-1}\mid g\in SL(2,\mathbb{R}),~{}v\in\mathbb{R}^{2}\}$.
-So, there are $g_{n}\in SL(2,\mathbb{R})$ and $v_{n}\in\mathbb{R}^{2}$ such that $(g_{n},v_{n})A(g_{n},v_{n})^{-1}\to H$.
-Using Iwasawa decomposition, we may assume that $g_{n}\in N^{+}$.
-Let $g_{n}=\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix}$ and $v_{n}=(\alpha_{n},\beta_{n})$.
-Suppose $(g,v)\in H$.
-Then, there are $a_{n}\in\mathbb{R}^{+}$ such that $(g_{n},v_{n})\mathfrak{a}_{n}(g_{n},v_{n})^{-1}\to(g,v)$, where $\mathfrak{a}_{n}=\begin{pmatrix}a_{n}&0\\ 0&1/a_{n}\end{pmatrix}$.
-This implies that
-$\begin{pmatrix}a_{n}&({1\over a_{n}}-a_{n})s_{n}\\ 0&{1\over a_{n}}\end{pmatrix}\to g,~{}~{}{\rm and}~{}~{}$
-$((1-a_{n})\alpha_{n}+(a_{n}-{1\over a_{n}})s_{n}\beta_{n},(1-{1\over a_{n}})\beta_{n})\to v.$
-By passing to subsequences and taking conjugates, we may assume that $s_{n}=0$ or $s_{n}\to\infty$, $\alpha_{n}=0$ or $\alpha_{n}\to\infty$ and $\beta_{n}=0$ or $\beta_{n}\to\infty$.
-We first consider the case that $s_{n}\to\infty$.
-If both $\alpha_{n}=0$ and $\beta_{n}=0$, then it is easy to see that $H=N^{+}=\left\{\begin{pmatrix}1&t\\ 0&1\end{pmatrix}\mid t\in\mathbb{R}\right\}$.
-So assume that either $\alpha_{n}\to\infty$ or $\beta_{n}\to\infty$.
-Then $v_{n}\to\infty$.
-This implies by Lemma 5.1 that $g$ is unipotent.
-Therefore, $a_{n}\to 1$.
-Since $g_{n}\in N^{+}$, and $B=AN^{+}$, it follows that $(g_{n},v_{n})A(g_{n},v_{n})^{-1}$ are closed subgroups of $B\ltimes\mathbb{R}^{2}$.
-Since $B\ltimes\mathbb{R}^{2}$ is isomorphic to a group of upper triangular matrices with positive diagonal entries, $B\ltimes\mathbb{R}^{2}$ is an exponential Lie group.
-Hence, by Proposition 2.2, $H$ is a connected one-dimensional subgroup of $N^{+}\ltimes\mathbb{R}^{2}$ which is the $3$-dimensional Heisenberg group.
-Therefore,
-$H=V_{(a,b,c)}=\left\{\left(\begin{pmatrix}1&at\\ 0&1\end{pmatrix},\left(ct+{abt^{2}\over 2},bt\right)\right)\mid t\in\mathbb{R}\right\}$
-for some $a,b,c\in\mathbb{R}$. Then for each $t\in\mathbb{R}$, there are $b_{n}$ (depending on $t$) such that
-$b_{n}\to 1,~{}~{}({1\over b_{n}}-b_{n})s_{n}\to at~{}~{}{\rm and}~{}~{}(1-{1\over b_{n}})\beta_{n}\to bt.$
-Case $1$: Suppose ${s_{n}\over\beta_{n}}\to r\in\mathbb{R}$. Then $a=-2br$ and hence
-$(1-b_{n})\alpha_{n}-{(1+b_{n})\over b_{n}}(1-b_{n})s_{n}\beta_{n}\to ct-rb^{2}t^{2}.$
-If $b=0$, then $H=V_{0}$. So, we may assume that $b\not=0$. Also
-$2(1-b_{n})s_{n}\beta_{n}-{(1+b_{n})\over b_{n}}(1-b_{n})s_{n}\beta_{n}\to-rb^{2}t^{2}.$
-Therefore,
-$(1-b_{n})\alpha_{n}-2(1-b_{n})s_{n}\beta_{n}\to ct.$
-This implies that $2s_{n}-{\alpha_{n}\over\beta_{n}}\to{c\over b}$. Let $d=\lim 2s_{n}-{\alpha_{n}\over\beta_{n}}$. Then $c=db$. Thus, $H=V_{(r_{1}b,r_{2}b,b)}$ for some $r_{1},r_{2}\in\mathbb{R}$.
-Case $2$: Suppose ${\beta_{n}\over s_{n}}\to p\in\mathbb{R}$, Then $b=-{pa\over 2}$ and hence
-$(1-b_{n})\alpha_{n}-{(1+b_{n})\over b_{n}}(1-b_{n})s_{n}\beta_{n}\to ct-{pa^{2}t^{2}\over 4}.$
-If $a=0$, then $H=V_{0}$. So, we may assume that $a\not=0$. Thus,
-$(1-b_{n})\alpha_{n}-2(1-b_{n})s_{n}\beta_{n}\to ct.$
-This implies that ${\alpha_{n}\over s_{n}}-2\beta_{n}\to 2{c\over a}$, hence $c={da\over 2}$ where $d=\lim{\alpha_{n}\over s_{n}}-2\beta_{n}$.
-Thus, $H=V_{(a,r_{1}a,r_{2}a)}$ for some $r_{1},r_{2}\in\mathbb{R}$.
-We now consider the case $s_{n}=0$. Then $g=I$ and
-$((1-a_{n})\alpha_{n},(1-{1\over a_{n}})\beta_{n})\to v$
-for some $a_{n}\to 1$.
-Case $3$: Suppose ${\beta_{n}\over\alpha_{n}}\to 0$.
-Then since $v_{n}=(\alpha_{n},\beta_{n})\to\infty$, $\alpha_{n}\to\infty$ and $(1-{1\over a_{n}})\beta_{n}={-1\over a_{n}}(1-a_{n})\alpha_{n}{\beta_{n}\over\alpha_{n}}\to 0$.
-Therefore $H\subset V_{0}$.
-For any $t\in\mathbb{R}$, taking $b_{n}=1-{t\over\alpha_{n}}$, we get that $b_{n}\to 1$, $(1-b_{n})\alpha_{n}=t$.
-Therefore, $((1-b_{n})\alpha_{n},(1-{1\over b_{n}})\beta_{n})\to(t,0)$.
-Thus, $H=V_{0}$.
-Case $4$: Suppose ${\alpha_{n}\over\beta_{n}}\to 0$.
-Then $(1-a_{n})\alpha_{n}=-a_{n}(1-{1\over a_{n}})\beta_{n}{\alpha_{n}\over\beta_{n}}\to 0$.
-Therefore, $H\subset V_{\infty}=\{(0,t)\mid t\in\mathbb{R}\}$.
-For any $t\in\mathbb{R}$, take $b_{n}={\beta_{n}\over\beta_{n}-t}$.
-Then $b_{n}\to 1$ and $((1-b_{n})\alpha_{n},(1-{1\over b_{n}})\beta_{n})\to(0,t)$.
-Thus, $H=V_{\infty}$.
-Case $5$: Suppose ${\beta_{n}\over\alpha_{n}}\to c\not=0$.
-Let $x,y\in\mathbb{R}$ be such that $v=(x,y)$.
-Then $(1-a_{n})\alpha_{n}\to x$ and $(1-{1\over a_{n}})\beta_{n}\to y$.
-Since $(1-{1\over a_{n}})\beta_{n}={-1\over a_{n}}(1-a_{n})\alpha_{n}{\beta_{n}\over\alpha_{n}}$, we get that $y=-cx$.
-Therefore $H\subset V_{r}=\{(t,rt)\mid t\in\mathbb{R}\}$ for $r=-c$.
-For $t\in\mathbb{R}$, let $b_{n}=1-{t\over\alpha_{n}}$.
-Then $b_{n}\to 1$ and $(1-b_{n})\alpha_{n}=t$.
-Also, $(1-{1\over b_{n}})\beta_{n}={-t\over\alpha_{n}-t}\beta_{n}\to-tc$.
-Therefore $V_{r}\subset H$.
-Thus, $H=V_{r}$.
-5.4. Limits of unipotent and Borel subgroups
-We now find the conjugacy limits of the Borel subgroup
-$B:=\left\{\left(\begin{pmatrix}a&s\\ 0&1/a\end{pmatrix}:\ a,s\in\mathbb{R}\right)\right\}$
-of $\operatorname{SL}(2,\mathbb{R})$ inside $G=\operatorname{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$.
-Proof of Theorem 1.10.
-Clearly, a conjugacy limit of $B$, when conjugated by elements of the type $(g,0)$ with $g\in\operatorname{SL}(2,\mathbb{R})$, is $B$ itself or a conjugate of $B$ by elements of $K$.
-For conjugation by more general type of elements, it is enough to consider conjugation by elements of the form $(I,v)$ with $v\in\mathbb{R}^{2}$.
-So, let $\mathfrak{g}_{n}=(I,v_{n})$ and $\mathfrak{b}_{n}=\begin{pmatrix}a_{n}&s_{n}\\ 0&1/a_{n}\end{pmatrix}$.
-Then a simple calculation shows that
-$\mathfrak{g}_{n}\mathfrak{b}_{n}\mathfrak{g}_{n}^{-1}=\left(\begin{pmatrix}a_{n}&s_{n}\\ 0&1/a_{n}\end{pmatrix},\Big{(}(1-a_{n})\alpha_{n}-s_{n}\beta_{n},\ (1-1/a_{n})\beta_{n}\Big{)}\right)$
-$=\big{(}\mathfrak{b}_{n},(x_{n},y_{n})\big{)},\ (\text{say})$
-where $v_{n}=(\alpha_{n},\beta_{n})$.
-Assume that $v_{n}\to\infty$ and $\mathfrak{g}_{n}\mathfrak{b}_{n}\mathfrak{g}_{n}^{-1}\to(g,v)$.
-Then by Lemma 5.1, $g$ is unipotent, hence $a_{n}\rightarrow 1$ necessarily.
-We shall consider the following cases.
-Case $1$: Suppose $\frac{\alpha_{n}}{\beta_{n}}\rightarrow 0$.
-Then $\beta_{n}\to\infty$.
-If $s_{n}\rightarrow s$ for some non-zero $s$, then it is easy to see that $x_{n}$ cannot converge.
-Hence, $s_{n}\rightarrow 0$.
-Given $x,y\in\mathbb{R}$, choose $s_{n}\rightarrow 0$ and $a_{n}\rightarrow 1$ in such a way that $-\beta_{n}s_{n}\rightarrow x$ and $(1-1/a_{n})\beta_{n}\rightarrow y$.
-Then $(1-a_{n})\alpha_{n}=-a_{n}(1-{1\over a_{n}})\beta_{n}{\alpha_{n}\over\beta_{n}}\to 0$, hence $\mathfrak{g}_{n}\mathfrak{b}_{n}^{\prime}\mathfrak{g}_{n}^{-1}\rightarrow\big{(}I,(x,y)\big{)}$ for suitable $\mathfrak{b}_{n}^{\prime}\in B$.
-This shows that, in this case, the limit subgroup is $\mathbb{R}^{2}$.
-Case $2$: Suppose $\frac{\beta_{n}}{\alpha_{n}}\rightarrow 0$.
-Then $\alpha_{n}\to\infty$.
-Let $x\in\mathbb{R}$.
-Then for $\tilde{\mathfrak{b}}_{n}=\begin{pmatrix}1-{x\over\alpha_{n}}&0\\ 0&{\alpha_{n}\over\alpha_{n}-x}\end{pmatrix}$, we get that $\mathfrak{g}_{n}\tilde{\mathfrak{b}_{n}}\mathfrak{g}_{n}^{-1}\to(I,(t,0))$.
-Thus, $\{(I,(t,0))\mid t\in\mathbb{R}\}\subset H$.
-If $\beta_{n}$ converges, then $y_{n}=(1-{1\over a_{n}})\beta_{n}\to 0$ and it can be shown that $H=N^{+}\ltimes\{(t,0)\mid t\in\mathbb{R}\}$.
-So, we may assume that $\beta_{n}\to\infty$.
-Since $\mathfrak{b}_{n}$ converges, $s_{n}$ converges, say to $s$.
-Since $\beta_{n}\to\infty$ and $x_{n}$ converges, we get that $(1-a_{n}){\alpha_{n}\over\beta_{n}}\to s$.
-If ${\beta_{n}^{2}\over\alpha_{n}}\to c\in\mathbb{R}$, then $y_{n}={-1\over a_{n}}(1-a_{n}){\alpha_{n}\over\beta_{n}}{\beta_{n}^{2}\over\alpha_{n}}\to-sc$.
-Therefore, by Proposition 2.2 we get that $H=\{(\mathfrak{s},(t,-cs)\mid t,s\in\mathbb{R}\}$ where $\mathfrak{s}=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}$.
-If ${\beta_{n}^{2}\over\alpha_{n}}\to\infty$, then $(1-a_{n}){\alpha_{n}\over\beta_{n}}=(1-{1\over a_{n}})\beta_{n}(-a_{n}){\alpha_{n}\over\beta_{n}^{2}}\to 0$.
-Therefore, by Proposition 2.2 we get that $H=\mathbb{R}^{2}=\{(I,(t,r)\mid t,r\in\mathbb{R}\}$.
-Case $3$: Suppose $\frac{\alpha_{n}}{\beta_{n}}\rightarrow c$ for some non-zero real number $c$.
-Note that we are necessarily assuming both $|\alpha_{n}|$ and $|\beta_{n}|$ are unbounded.
-Otherwise, the limit subgroup will be a conjugate of $B$.
-Now, suppose $x_{n}\rightarrow x$.
-Then $\frac{x_{n}}{\beta_{n}}\rightarrow 0$.
-But,
-$\frac{x_{n}}{\beta_{n}}=\frac{(1-a_{n})\alpha_{n}}{\beta_{n}}-s_{n}$
-which implies that $s_{n}\rightarrow 0$ in this case as well. Therefore $H\subset\mathbb{R}^{2}$. Thus, by Proposition 2.2, $H$ is of dimension $2$ and hence $H=\mathbb{R}^{2}$.
-We next determine the conjugacy limits of the unipotent subgroup $N^{+}$.
-Proof.
-Proof of Theorem 1.11 Since $N^{+}$ is normal in $B$, using Iwasawa decomposition in $SL(2,\mathbb{R})$, it follows that whenever $g_{n}\in\operatorname{SL}(2,\mathbb{R})$, the limit of $g_{n}N^{+}g_{n}^{-1}$ is either $N^{+}$ or its conjugates by elements of $K$.
-For conjugation by other elements of $G$, without loss of generality, we may consider conjugation by elements of the form $\mathfrak{g}_{n}=(k_{n},v_{n})$ with $k_{n}\in K$ and $v_{n}=(\alpha_{n},\beta_{n})\in\mathbb{R}^{2}$.
-Also, it is easy to see that if $L$ is a closed subgroup of $G$, $g_{n}Lg_{n}^{-1}\rightarrow H$ and $k_{n}\rightarrow k$, then $k_{n}g_{n}Lg_{n}^{-1}k_{n}^{-1}\rightarrow kHk^{-1}$.
-In view of this, it is enough to consider conjugation of $N^{+}$ by elements of the form $\mathfrak{g}_{n}=(I,v_{n})$.
-Now, suppose $\mathfrak{g}_{n}=\big{(}I,v_{n}\big{)}$ with $v_{n}=(\alpha_{n},\beta_{n})$. Then for any $\mathfrak{s}_{n}=\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix}$,
-$\mathfrak{g}_{n}\mathfrak{s}_{n}\mathfrak{g}_{n}^{-1}=\left(\begin{pmatrix}1&s_{n}\\ 0&1\end{pmatrix},\big{(}-s_{n}\beta_{n},0\big{)}\right).$
-If $|\beta_{n}|\rightarrow\infty$, then choosing $s_{n}\rightarrow 0$ suitably $\big{(}I,(t,0)\big{)}$ can be realized as a limit of $\mathfrak{g}_{n}\mathfrak{s}_{n}\mathfrak{g}_{n}^{-1}$ for any real number $t$.
-Also, if $\mathfrak{g}_{n}\mathfrak{s}_{n}\mathfrak{g}_{n}^{-1}$ converges, then $s_{n}$ as well as $s_{n}\beta_{n}$ converges.
-but, as $|\beta_{n}|\to\infty$, we have $s_{n}\to 0$.
-This shows that, in this case, the limit of $\mathfrak{g}_{n}N^{+}\mathfrak{g}_{n}^{-1}$ is $V_{0}$.
 Acknowledgement .
-The first author thanks the Indian Satitistical Institute Bangalore Center for its hospitality during his several visits while this work was going on.
-The second author acknowledges the support from SERC through the grant under MATRICS MTR/2022/000429.
+Both the authors would like to thank the anonymous referee who suggested the splitting of the previous version of the article, and many other valuable suggestions on improving the exposure of the article especially suggesting [1].
+The authors also like to thank International Centre for Theoretical Sciences, Bangalore for its hospitality during the programme Ergodic Theory and Dynamical Systems, $5$th December to $16$th December, 2022 when this work was initiated.
+The first author thanks the Indian Statistical Institute Bangalore Centre for its hospitality during his several visits while this work was going on.
+The second author acknowledges the support from SERB through the grant under MATRICS MTR/2022/000429.